### v2

# HiTeC: Hierarchical Contrastive Learning on Text-Attributed Hypergraph with Semantic-Aware Augmentation

Abstract
Contrastive learning (CL) has become a dominant paradigm for self-supervised hypergraph learning, enabling effective training without costly labels. However, node entities in real-world hypergraphs are often associated with rich textual information, which is overlooked in prior works. Directly applying existing CL-based methods to such text-attributed hypergraphs (TAHGs) leads to three key limitations: (1) The common use of graph-agnostic text encoders overlooks the correlations between textual content and hypergraph topology, resulting in suboptimal representations. (2) Their reliance on random data augmentations introduces noise and weakens the contrastive objective. (3) The primary focus on node- and hyperedge-level contrastive signals limits the ability to capture long-range dependencies, which is essential for expressive representation learning. Although HyperBERT pioneers CL on TAHGs, its co-training paradigm suffers from poor scalability. To fill the research gap, we introduce HiTeC, a two-stage hierarchical contrastive learning framework with semantic-aware augmentation for scalable and effective self-supervised learning on TAHGs. In the first stage, we pre-train the text encoder with a structure-aware contrastive objective to overcome the graph-agnostic nature of conventional methods. In the second stage, we introduce two semantic-aware augmentation strategies, including prompt-enhanced text augmentation and semantic-aware hyperedge drop, to facilitate informative view generation. Furthermore, we propose a multi-scale contrastive loss that extends existing objectives with an $s$-walk-based subgraph-level contrast to better capture long-range dependencies. By decoupling text encoder pretraining from hypergraph contrastive learning, this two-stage design enhances scalability without compromising representation quality. Extensive experiments on six real-world datasets validate the effectiveness of our proposed method.
1 Introduction
Hypergraphs extend traditional graphs by allowing each hyperedge to connect multiple nodes, offering a natural way to model multi-way relationships. Such higher-order structures are prevalent in complex real-world systems, e.g., cortical co-activation in brains (Yu et al. 2011), user-item interactions in e-commerce platforms (Xia, Huang, and Zhang 2022), and multi-author collaborations in academic networks (Kim et al. 2023). In these scenarios, nodes are often associated with rich textual attributes (e.g., product reviews in user–item networks and paper abstracts in co-authorship networks), which provide essential semantic information for downstream applications. While Hypergraph Neural Networks (HNNs) have emerged as a powerful paradigm for hypergraph mining (Chien et al. 2022b; Kim et al. 2024b), effectively capturing high-order structural dependencies, they primarily rely on numerical features and task-specific labels. This reliance limits their effectiveness in real-world applications, where labeled data is scarce and node attributes are predominantly textual. Therefore, underscoring the necessity and significance of self-supervised learning (SSL) on Text-Attributed Hypergraphs (TAHGs).
Recent advances in hypergraph self-supervised learning (HSSL) have primarily focused on contrastive learning paradigms, which aim to maximize the agreement between two augmented views derived from the original hypergraph. HyperGCL (Wei et al. 2022) pioneers this line of work by employing a learnable augmentation strategy. TriCL (Lee and Shin 2023) introduces tri-directional contrastive objectives that model node-, hyperedge-, and membership-level relations in hypergraphs. SE-HSSL (Li et al. 2024) extends this framework by re-designing a hierarchical membership-level objective, achieving state-of-the-art (SOTA) performance. Beyond contrastive methods, generative approaches have also gained initial attention. HypeBoy (Kim et al. 2024a) proposes a hyperedge-filling pretext task to learn expressive hypergraph representations.
Despite recent progress, current HSSL approaches still encounter significant limitations when adopted to hypergraphs with textual attributes: (1) They typically use shallow encoders (e.g., Bag-of-Words (Zhang, Jin, and Zhou 2010) and Skip-Gram (Mikolov et al. 2013)) or pre-trained language models (e.g., BERT (Devlin et al. 2019)) to convert raw text into numerical features, ignoring hypergraph structure. This graph-agnostic design overlooks the correlation between text and higher-order topology, leading to suboptimal representations (Chien et al. 2022a). (2) Current hypergraph contrastive learning (HCL) methods often apply random augmentations to features and structure, such as randomly masking text or hyperlinks. Such perturbations can distort textual meaning and disrupt the alignment between structure and semantic content, thereby introducing noise into the contrastive objective (Fang et al. 2024). (3) Existing HCL methods primarily focus on node- and hyperedge-level signals, while overlooking long-range dependencies and complex topological contexts, which are essential for learning expressive representations (Zhang et al. 2024).
HyperBERT (Bazaga, Lio, and Micklem 2024) takes an early attempt at SSL on TAHGs by jointly encoding textual semantics and hypergraph structure via a Transformer-based architecture. However, its co-training paradigm incurs substantial memory overhead, limiting scalability even on medium-sized hypergraphs (e.g., History and Photo (Yan et al. 2023)), as demonstrated in our experiments. Moreover, it focuses solely on node-level contrastive signals, ignoring multi-scale structural information that has proven beneficial for HSSL (Lee and Shin 2023; Li et al. 2024). Although recent text-attributed graph studies (Chien et al. 2022a; Fang et al. 2024; Zhang et al. 2024) explore joint modeling of text and structure, they overlook high-order relations in hypergraphs, which may limit their applicability to TAHGs.
To address these issues, we propose HiTeC, a hierarchical contrastive learning framework with semantic-aware augmentation for scalable and effective SSL on TAHGs. HiTeC adopts a two-stage design that decouples the pretraining of the text and hypergraph encoders, significantly improving the scalability over joint training paradigms. To overcome the graph-agnostic nature of existing text encoders, we first introduce a structure-aware contrastive objective during text encoder pretraining. Positive and negative text pairs are constructed based on structural proximity, encouraging the encoder to capture correlations between textual content and higher-order topology. Furthermore, to mitigate the noise introduced by random augmentations, we propose two semantic-aware augmentation strategies: a prompt-enhanced text augmentation module, which injects structural context into raw text to generate diverse yet semantically consistent views, and a semantic-aware hyperedge drop mechanism, which discards hyperedges according to their internal semantic cohesion to preserve meaningful relational patterns. Finally, to capture long-range dependencies and multi-scale structural knowledge, we formulate hierarchical contrastive objectives at the node, hyperedge, and subgraph levels. Specifically, our subgraph contrastive objective leverages an $s$-walk-based sampling strategy to extract structurally informative subgraphs, facilitating representation learning over broader contexts. Our main contributions are summarized as follows:
  • We propose HiTeC, a scalable and effective two-stage contrastive learning framework for TAHGs, addressing key limitations of existing HSSL methods in modeling complex structural dependencies and rich textual content.
  • We introduce a structure-aware contrastive objective to pre-train the text encoder, enabling integration of textual content with hypergraph structure, and overcoming the graph-agnostic nature of existing text encoders.
  • We enhance the hypergraph encoder with semantic-aware augmentations and hierarchical contrastive objectives. Specifically, a prompt-enhanced text augmentation and semantic-aware hyperedge drop strategies are designed to reduce perturbation noise. And a novel $s$-walk-based subgraph-level signal is employed to capture long-range structural dependencies.
  • We conduct a comprehensive evaluation of HiTeC on six real-world datasets. The experimental results demonstrate its superiority over strong baselines across various benchmark downstream tasks.
2 Preliminary
Notations. We define a text-attributed hypergraph (TAHG) as $\mathcal{H}=\{\mathcal{V},\mathcal{E},\mathcal{T}\}$, where $\mathcal{V}=\{v_{i}\}_{i=1}^{|\mathcal{V}|}$ denotes the set of nodes, and $\mathcal{E}=\{e_{j}\}_{j=1}^{|\mathcal{E}|}$ denotes the set of hyperedges, with each hyperedge $e\in\mathcal{E}$ is a non-empty subset of nodes, and $|e|$ is the size of $e$. Each node $v_{i}$ is associated with a textual sequence $t_{i}\in\mathcal{T}$, and $\mathcal{T}$ represents the set of all text attributes. Each hyperedge $e_{j}$ is associated with a positive scalar weight $w_{j}$, and all weights formulate a diagonal matrix $\mathbf{W}\in\mathbb{R}^{|\mathcal{E}|\times|\mathcal{E}|}$. The hypergraph structure can be modeled as an incidence matrix $\mathbf{H}\in\{0,1\}^{|\mathcal{V}|\times|\mathcal{E}|}$, where $h_{ij}=1$ if node $v_{i}$ belongs to hyperedge $e_{j}$, and $h_{ij}=0$ otherwise. We define the node degree matrix $\mathbf{D}_{v}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$, where each entry is $d(v_{i})=\sum_{e_{j}\in\mathcal{E}}w_{j}\cdot h_{ij}$. Similarly, the hyperedge degree matrix is defined as $\mathbf{D}_{e}\in\mathbb{R}^{|\mathcal{E}|\times|\mathcal{E}|}$, where each element $\delta(e_{j})=\sum_{v_{i}\in e_{j}}h_{ij}$ represents the number of nodes connected by the hyperedge $e_{j}$.
Hypergraph neural networks. Hypergraph neural networks (HNNs) have emerged as effective models for representation learning on hypergraphs (Feng et al. 2019; Bai, Zhang, and Torr 2021; Chien et al. 2022b). Most HNNs adopt a two-stage message passing mechanism that alternates between node-to-hyperedge and hyperedge-to-node aggregation. At each layer $l$, the embedding of a hyperedge is updated based on its incident nodes, while a node embedding is updated via its associated hyperedges:
$\begin{split}z^{(l)}_{e_{j}}&=f^{(l)}_{\mathcal{V}\rightarrow\mathcal{E}}\left(z^{(l-1)}_{e_{j}},\{z^{(l-1)}_{v_{k}}\mid v_{k}\in e_{j}\}\right),\\ z^{(l)}_{v_{i}}&=f^{(l)}_{\mathcal{E}\rightarrow\mathcal{V}}\left(z^{(l-1)}_{v_{i}},\{z^{(l)}_{e_{k}}\mid v_{i}\in e_{k}\}\right),\end{split}$ (1)
where $z^{(l)}_{e_{j}}$ and $z^{(l)}_{v_{i}}$ denote the embeddings of hyperedge $e_{j}$ and node $v_{i}$ at layer $l$, respectively. The functions $f^{(l)}_{\mathcal{V}\rightarrow\mathcal{E}}$ and $f^{(l)}_{\mathcal{E}\rightarrow\mathcal{V}}$ are permutation-invariant aggregation functions that propagate messages between nodes and hyperedges. This iterative process allows the model to capture complex high-order relations in the hypergraph structure.
Problem statement. Given a TAHG $\mathcal{H}=\{\mathcal{V},\mathcal{E},\mathcal{T}\}$, the objective is to learn an optimal function $\Psi:\mathcal{H}\rightarrow(\mathbf{Z}_{\mathcal{V}},\mathbf{Z}_{\mathcal{E}})$ in an unsupervised manner that maps the hypergraph to low-dimensional embeddings of nodes and hyperedges. These representations may then be applied to downstream tasks such as node classification and hyperedge prediction.
3 Methodology
Figure 1: The overall framework of HiTeC.
3.1 Overview
As illustrated in Figure 1, we present HiTeC, a contrastive learning framework tailored for TAHGs. In the first stage, we pre-train the text encoder $f_{\theta}$ with a structure-aware contrastive objective, allowing it to capture topological knowledge and achieve more expressive text representations than graph-agnostic encoders. In the second stage, we pre-train a hypegraph encoder $g_{\phi}$ with $f_{\theta}$ frozen. To fully exploit the rich semantic context of nodes, we introduce two semantic-aware augmentation strategies on both text- and structure-level, enhancing view diversity while preserving semantic coherence. Furthermore, we design a hierarchical contrastive objective across node, hyperedge, and subgraph levels. Specifically, the subgraph-level objective, enabled by $s$-walk sampling, captures long-range dependencies. By decoupling the training of two encoders, HiTeC achieves better scalability and performance in integrating textual content and high-order structures.
3.2 Structure-aware Text Encoder Pretraining
Text encoder. Our text encoder $f_{\theta}$ is built upon a pre-trained language model (PLM) (e.g., BERT (Devlin et al. 2019)) with a lightweight projection layer. To enable efficient training, we apply the parameter-efficient approach, LoRA (Hu et al. 2022). Given a node $v_{i}\in\mathcal{V}$ with textual attribute $t_{i}$, the textual representation is formulated as:
$x_{i}=f_{\theta}(t_{i}),$ (2)
where $x_{i}\in\mathbb{R}^{d}$ denotes the embedding corresponding to the [CLS] token in the output sequence.
Structure-aware contrastive objective. In real-world hypergraphs, nodes are often semantically closer to their local neighbors than to distant ones. For example, in co-citation hypergraphs, papers within the same hyperedge tend to be more topically related. To encode these structural priors, we pre-train $f_{\theta}$ by contrasting local and distant semantic contexts. Specifically, for each node $v_{i}\in\mathcal{V}$, we treat its 1-hop neighbors $N^{(1)}_{v_{i}}$ as positives, and nodes beyond 1-hop distance $\mathcal{V}\setminus N^{(1)}_{v_{i}}$ as negatives. As shown in Figure 1, 1-hop neighbors are defined as nodes belong to the same hyperedges that contain $v_{i}$. Contrastive pairs are constructed by mean pooling embeddings from each set:
$\begin{split}x_{i}^{+}&=\frac{1}{|N^{(1)}_{v_{i}}|}\sum_{v_{j}\in N^{(1)}_{v_{i}}}f_{\theta}(t_{j}),\\ x_{i}^{-}&=\frac{1}{|\mathcal{V}\setminus N^{(1)}_{v_{i}}|}\sum_{v_{j}\in\mathcal{V}\setminus N^{(1)}_{v_{i}}}f_{\theta}(t_{j}),\end{split}$ (3)
where $x_{i}^{+}\in\mathbb{R}^{d}$ and $x_{i}^{-}\in\mathbb{R}^{d}$ denote the aggregated representations of the positive and negative samples, respectively.
A triplet margin loss (Schroff, Kalenichenko, and Philbin 2015) is adopted to guide $f_{\theta}$ in distinguishing positive sample $x_{i}^{+}$ from negative sample $x_{i}^{-}$. For each anchor $x_{i}$, the loss of the text encoder is defined as:
$\ell_{t}(x_{i},x_{i}^{+},x_{i}^{-})=\max\{\cos(x_{i},x_{i}^{+})-\cos(x_{i},x_{i}^{-})+m\},$ (4)
where $\cos(\cdot)$ denotes cosine similarity, and $m$ is the margin hyperparameter. The overall structure-aware objective is averaged across all nodes:
$\mathcal{L}_{t}=\frac{1}{|\mathcal{V}|}\sum_{v_{i}\in\mathcal{V}}\ell_{t}(x_{i},x_{i}^{+},x_{i}^{-}),$ (5)
This process allows $f_{\theta}$ to incorporate structural signals into textual representations.
Figure 2: An example of the prompt template on Cora.
3.3 Semantic-aware Hypergraph Encoder Pretraining
Existing HCL methods often overlook node textual semantics and rely on random augmentation strategies, resulting in suboptimal performance (Fang et al. 2024; Wu et al. 2024). Moreover, they struggle to model long-range dependencies, limiting broader contextual understanding (Zhang et al. 2024; Wu et al. 2024). To tackle these issues, we propose a semantic-aware contrastive learning strategy to pre-train $g_{\phi}$. As shown in Figure 1, this stage comprises three components: semantic-aware augmentation, hypergraph encoding, and hierarchical contrastive objectives.
Semantic-aware augmentation. To mitigate the noise from random perturbations, we introduce two novel augmentation techniques: prompt-enhanced text augmentation, which generates diverse yet semantically consistent textual views, and semantic-aware hyperedge drop, which ensures structural perturbations remain aligned with the semantics.
Prompt-enhanced text augmentation. As feature-level perturbations can easily distort semantic meaning, we instead focus on augmenting the raw text using a prompt-enhanced strategy. For each node text $t_{i}$, we construct an augmented view $\tilde{t}_{1,i}=\text{Prompt}(t_{i},\mathcal{H})$, while retaining the original text as $\tilde{t}_{2,i}=t_{i}$. As shown in Figure 2,the prompt template integrates the original text with domain knowledge, global topology, and neighbor context, concatenated into a single sequence. Each component adds a distinct perspective: domain knowledge offers semantic background, global topology introduces structural awareness, and neighbor context provides complementary local semantics. This simple yet effective strategy generates semantically consistent but contextually diverse text variants, with a shared $f_{\theta}$ that maps them to a unified space, denoted as $\tilde{\mathbf{X}}_{1}$ and $\tilde{\mathbf{X}}_{2}$, respectively.
Semantic-aware hyperedge drop. Effective structure augmentation should consider both textual semantics and the original topology. To achieve that, we define a semantic cohesiveness score that reflects both the structural grouping of nodes and their semantic similarity. This score is then used to guide the drop probability of each node–hyperedge incidence: connections within more cohesive hyperedges are more likely to be preserved. In this way, our strategy encourages the retention of structurally and semantically aligned patterns, while selectively pruning noisy or weakly related hyperlinks. We formally define the it as follows.
Definition 1 (semantic cohesiveness score).
Given a TAHG $\mathcal{H}(\mathcal{V},\mathcal{E},\mathbf{X})$, the semantic cohesiveness score of a hyperedge $e\in\mathcal{E}$ is defined as the average pairwise cosine similarity among its member nodes. Let $\mathbf{X}_{e}\in\mathbb{R}^{|e|\times d}$ denote the normalized feature matrix for nodes in $e$. The semantic cohesiveness score $s(e)$ is then computed as:
$s(e)=\frac{1}{\binom{|e|}{2}}\sum_{i=1}^{|e|}\sum_{j=i+1}^{|e|}\bm{S}_{e}(i,j),\quad\bm{S}_{e}=\mathbf{X}_{e}\mathbf{X}_{e}^{\top},$
where $\bm{S}_{e}(i,j)$ denotes the similarity between the $i$-th and $j$-th node in hyperedge $e$.
The drop probability of each node-hyperedge incidence $(v_{i},e_{j})$ where $v_{i}\in e_{j}$ is then guided by $s(e_{j})$:
$p_{\text{drop}}(v_{i},e_{j})=1-\sigma\left(\frac{s(e_{j})-0.5}{\tau_{\text{drop}}}\right),$ (6)
where $\sigma(\cdot)$ is the sigmoid function and $\tau_{\text{drop}}$ is a temperature hyperparameter. Higher cohesiveness scores lead to lower drop probabilities. In implementation, we construct a binary masking matrix $\mathbf{M}\in\{0,1\}^{|\mathcal{V}|\times|\mathcal{E}|}$, where each entry $m_{ij}\sim\mathcal{B}(1-p_{\text{drop}}(v_{i},e_{j}))$ is sampled from a Bernoulli distribution. The perturbed incidence matrix $\tilde{\mathbf{H}}$ is represented as:
$\tilde{\mathbf{H}}=\mathbf{M}\odot\mathbf{H},$ (7)
where $\odot$ denotes element-wise multiplication. We apply this semantic-aware hyperedge drop mechanism to both views, yielding two augmented hypergraphs: $\tilde{\mathcal{H}}_{1}=(\tilde{\mathbf{X}}_{1},\tilde{\mathbf{H}}_{1})$ and $\tilde{\mathcal{H}}_{2}=(\tilde{\mathbf{X}}_{2},\tilde{\mathbf{H}}_{2})$. This procedure complements text-level augmentation by introducing structural diversity while retaining semantic consistency, enabling the hypergraph encoder to learn more robust node representations.
Hypergraph encoder. We take the HGNN (Feng et al. 2019) with the element-wise mean pooling layer as our hypergraph encoder. At the $l$-th layer, the hyperedge representation $\mathbf{Z}_{\mathcal{E}}^{(l)}\in\mathbb{R}^{|\mathcal{E}|\times d}$ and the node representation $\mathbf{Z}_{\mathcal{V}}^{(l)}\in\mathbb{R}^{|\mathcal{V}|\times d}$ are updated as:
$\begin{split}\mathbf{Z}_{\mathcal{E}}^{(l)}&=\rho\left(\mathbf{D}_{e}^{-1}\mathbf{H}^{\top}\mathbf{Z}_{\mathcal{V}}^{(l-1)}\mathbf{\Theta}_{\mathcal{E}}^{(l)}+\bm{b}_{e}^{(l)}\right),\\ \mathbf{Z}_{\mathcal{V}}^{(l)}&=\rho\left(\mathbf{D}_{v}^{-1}\mathbf{H}\mathbf{W}\mathbf{Z}_{\mathcal{E}}^{(l)}\mathbf{\Theta}_{\mathcal{V}}^{(l)}+\bm{b}_{v}^{(l)}\right),\end{split}$ (8)
where $\rho(\cdot)$ is the PReLU activation function, $\mathbf{\Theta}_{\mathcal{E}}^{(l)}$ and $\mathbf{\Theta}_{\mathcal{V}}^{(l)}$ are trainable weight matrices, and $\bm{b}_{e}^{(l)}$ and $\bm{b}_{v}^{(l)}$ are corresponding bias terms. $\mathbf{W}$ is initialized as an identity matrix, which means equal weights for all hyperedges.
Hierarchical contrastive objectives. Prior works (Lee and Shin 2023; Li et al. 2024) have demonstrated that multi-scale contrastive objectives are beneficial for HSSL. To better capture structural semantics at different granularities, we adopt a hierarchical contrastive learning objective spanning node, hyperedge, and subgraph levels. In particular, the novel subgraph-level objective, based on $s$-walk sampling, enables the model to capture long-range dependencies beyond local structure. All objectives are optimized with InfoNCE loss (Oord, Li, and Vinyals 2018).
[…]
$\ell_{n}(\mathbf{z}_{1,v_{i}},\mathbf{z}_{2,v_{i}})=-\log\frac{e^{\cos(\mathbf{z}_{1,v_{i}},\mathbf{z}_{2,v_{i}})/\tau_{n}}}{\sum_{k=1}^{|\mathcal{V}|}e^{\cos(\mathbf{z}_{1,v_{i}},\mathbf{z}_{2,v_{k}})/\tau_{n}}},$ (9)
where $\tau_{n}$ is the temperature coefficient. In practice, we symmetrize the loss by using the second view’s embedding as the anchor and obtain the final node-level contrastive loss:
$\mathcal{L}_{n}=\frac{1}{2|\mathcal{V}|}\sum_{i=1}^{|\mathcal{V}|}\left[\ell_{n}(\mathbf{z}_{1,v_{i}},\mathbf{z}_{2,v_{i}})+\ell_{n}(\mathbf{z}_{2,v_{i}},\mathbf{z}_{1,v_{i}})\right].$ (10)
Hyperedge-level objective. To consider group-level knowledge, we integrate hyperedge-level contrastive signal during optimization. Given a hyperedge $e_{j}\in\mathcal{E}$, let $\mathbf{z}_{1,e_{j}}$ and $\mathbf{z}_{2,e_{j}}$ be its hyperedge representations from the two augmented views. Similar to the node-level objective, the contrastive loss with $\mathbf{z}_{1,e_{j}}$ as anchor is formulated as:
$\ell_{\text{e}}(\mathbf{z}_{1,e_{j}},\mathbf{z}_{2,e_{j}})=-\log\frac{e^{\cos(\mathbf{z}_{1,e_{j}},\mathbf{z}_{2,e_{j}})/\tau_{e}}}{\sum_{k=1}^{|\mathcal{E}|}e^{\cos(\mathbf{z}_{1,e_{j}},\mathbf{z}_{2,e_{k}})/\tau_{e}}},$ (11)
where $\tau_{e}$ is the temperature coefficient. Symmetrizing the loss, the final hyperedge-level objective can be written as:
$\mathcal{L}_{e}=\frac{1}{2|\mathcal{E}|}\sum_{j=1}^{|\mathcal{E}|}\left[\ell_{e}(\mathbf{z}_{1,e_{j}},\mathbf{z}_{2,e_{j}})+\ell_{e}(\mathbf{z}_{2,e_{j}},\mathbf{z}_{1,e_{j}})\right].$ (12)
Subgraph-level objective.
Figure 3: An example of the $s$-walk-based subgraph sampling process. For center node $v_{c}$, we perform $s$-walk with $s=2$ and $l=4$ on $\mathcal{H}$. The hyperedge sequences visited are $e_{1},e_{3},e_{4},e_{5}]$. The sampled subgraph is denoted as $\mathcal{H}_{s_{c}}$.
While node- and hyperedge-level objectives capture local semantic and structural signals, they may fail to model long-range dependencies in hypergraphs. To bridge this gap, we further introduce a subgraph-level objective by constructing broader structural signals through a novel $s$-walk-based sampling strategy. Unlike random walk, $s$-walk is a high-order sequence in hypergraphs, where $s$ measures the overlap strength between hyperedges, enabling the extraction of cohesive and semantically meaningful substructures unique to hypergraphs (Aksoy et al. 2020; Preti, De Francisci Morales, and Bonchi 2024). We next define $s$-walk and detail our subgraph sampling method.
Definition 2 ($s$-walk).
Given $s\in\mathbb{N}^{+}$, an $s$-walk of length $l$ is a sequence of hyperedges $e_{1},\ldots,e_{l}]$ such that each consecutive pair of hyperedges shares at least $s$ nodes, i.e., $|e_{j}\cap e_{j+1}|\geq s$ for all $j\in[1,l-1]$.
As illustrated in Figure 3, we construct an $s$-walk-based subgraph $\mathcal{H}_{s_{c}}=\{\mathcal{V}_{s_{c}},\mathcal{E}_{s_{c}}\}$ by first selecting a starting hyperedge $e_{1}$ that contains the center node $v_{c}$, and then performing an $s$-walk of length $l$. We randomly sample one node from each visited hyperedge. The sampled subgraph includes all visited hyperedges and their associated nodes: $\mathcal{E}_{s_{c}}=\{e_{j}\}_{j=1}^{l}$, $\mathcal{V}_{s_{c}}=\bigcup_{j=1}^{l}e_{j}$. Each subgraph $\mathcal{H}_{s_{c}}$ can be encoded by the shared hypergraph encoder $g_{\phi}(\cdot)$, and its representation is obtained via mean pooling $P_{\text{mean}}$ over the node embeddings:
$\mathbf{z}_{s_{c}}=P_{\text{mean}}\left(g_{\phi}(\mathcal{H}_{s_{c}})\right).$ (13)
However, constructing and encoding subgraphs for all nodes incurs a high computational cost. To mitigate this, we follow prior findings suggesting that high-degree nodes tend to carry richer contextual information (Fang et al. 2025). Specifically, we select a subset of representative nodes $\mathcal{V}_{r}\subseteq\mathcal{V}$ by ranking node degrees and retaining the top $r\%$. As validated in our experiments, this strategy maintains competitive performance even with a significantly reduced number of subgraphs. For each anchor node $v_{m}\in\mathcal{V}_{r}$, we obtain subgraph representations $\mathbf{z}_{1,s_{m}}$ and $\mathbf{z}_{2,s_{m}}$ from two augmented views. Similar to the former contrastive signals, the subgraph-level loss with anchor $\mathbf{z}_{1,s_{m}}$ is defined as:
$\ell_{\text{s}}(\mathbf{z}_{1,s_{m}},\mathbf{z}_{2,s_{m}})=-\log\frac{e^{\cos(\mathbf{z}_{1,s_{m}},\mathbf{z}_{2,s_{m}})/\tau_{s}}}{\sum_{k=1}^{|\mathcal{V}_{r}|}e^{\cos(\mathbf{z}_{1,s_{m}},\mathbf{z}_{2,s_{k}})/\tau_{s}}},$ (14)
where $\tau_{s}$ is a temperature coefficient. And the final symmetric loss across all anchors is:
$\mathcal{L}_{\text{s}}=\frac{1}{2|\mathcal{V}_{r}|}\sum_{m=1}^{|\mathcal{V}_{r}|}\left[\ell_{s}(\mathbf{z}_{1,s_{m}},\mathbf{z}_{2,s_{m}})+\ell_{\text{s}}(\mathbf{z}_{2,s_{m}},\mathbf{z}_{1,s_{m}})\right].$ (15)
Overall objective. Finally, by integrating Eq. (10), (12), and (15), our proposed contrastive loss is formulated as:
$\mathcal{L}=\mathcal{L}_{n}+\lambda_{e}\mathcal{L}_{e}+\lambda_{s}\mathcal{L}_{s},$ (16)
where $\lambda_{e}$ and $\lambda_{s}$ are balancing weights. This hierarchical contrastive framework enables the model to capture patterns at multiple granularities, generating more robust and discriminative node representations.
4 Experiments
Dataset $|\mathcal{V}|$ $|\mathcal{E}|$ Avg.|e||e| # Avg.tokens # Class Citeseer 1,778 2,118 2 198 6 Cora 2,708 1,579 3 189 7 History 41,551 169,454 9 300 12 Photo 48,362 212,247 11 189 12 Computers 87,229 277,539 9 115 10 Fitness 173,055 1,468,229 11 28 13
Table 1: Statistics of datasets.
Method Citeseer Cora History Photo Computers Fitness
Graph SSL
GraphCL + SE 59.63 ± 1.58 56.69 ± 1.99 62.86 ± 0.22 42.61 ± 0.21 34.45 ± 0.38 52.11 ± 0.43
GraphCL + GIANT 57.58 ± 1.90 55.64 ± 1.50 62.08 ± 0.24 42.74 ± 0.22 34.65 + 0.23 52.04 ± 0.34
BGRL + SE 60.81 ± 1.45 59.72 ± 1.32 65.74 ± 0.30 44.60 ± 0.30 39.40 ± 0.46 55.66 ± 0.33
BGRL + GIANT 58.68 ± 1.92 57.44 ± 1.39 63.99 ± 0.35 43.65 ± 0.25 37.00 ± 0.37 55.05 ± 0.25
GraphMAE + SE 46.42 ± 2.75 50.33 ± 0.47 52.90 ± 9.01 41.52 ± 0.08 25.61 ± 0.11 40.65 ± 0.37
GraphMAE + GIANT 50.11 ± 1.74 53.59 ± 1.07 56.20 ± 0.12 41.54 ± 0.06 25.57 ± 0.12 42.04 ± 0.04
HSSL
TriCL + SE 60.78 ± 1.00 58.79 ± 1.30 69.28 ± 0.32 47.11 ± 0.44 43.43 ± 0.66 45.46 ± 0.21
TriCL + BERT 64.50 ± 1.08 62.67 ± 1.13 75.01 ± 0.22 51.25 ± 0.62 46.60 ± 0.90 48.97 ± 0.63
TriCL + RoBERTa 64.48 ± 0.95 60.74 ± 1.01 70.62 ± 0.81 48.42 ± 0.67 43.82 ± 1.11 47.49 ± 0.60
SE-HSSL + SE 57.04 ± 3.31 59.87 ± 1.83 68.59 ± 0.74 43.75 ± 2.01 36.53 ± 2.32 54.67 ± 2.00
SE-HSSL + BERT 65.83 ± 1.21 64.30 ± 0.89 74.88 ± 0.42 51.80 ± 0.63 41.63 ± 2.21 56.92 ± 1.44
SE-HSSL + RoBERTa 65.04 ± 1.04 62.76 ± 1.15 72.13 ± 1.07 49.41 ± 2.62 40.86 ± 3.99 54.79 ± 0.64
HypeBoy + SE 59.98 ± 2.45 29.87 ± 0.62 OOM OOM OOM OOM
HypeBoy + BERT 61.98 ± 1.70 32.73 ± 1.71 OOM OOM OOM OOM
HypeBoy + RoBERTa 50.41 ± 2.88 30.84 ± 0.95 OOM OOM OOM OOM
VilLain (w/o feat.) 42.22 ± 1.72 53.72 ± 1.00 40.99 ± 3.06 33.94 ± 1.39 20.99 ± 0.91 OOM
HyperBERT 30.25 ± 3.04 26.85 ± 0.66 OOM OOM OOM OOM
HiTeC 66.38 ± 1.73 74.07 ± 0.88 79.81 ± 0.16 59.64 ± 0.16 55.81 ± 0.12 67.18 ± 0.19
Table 2: Node classification accuracy (% mean ± std) across six datasets. The best and second-best results are highlighted in bold and underline, respectively. “w/o feat.” means without node features. OOM denotes out-of-memory on a 24GB GPU.
4.1 Experimental Setup
Datasets. Since there are no publicly available TAHG datasets, we introduce six new text-attributed hypergraph datasets using a max clique-based hypergraph reconstruction method (Bron and Kerbosch 1973; Wang and Kleinberg 2024). These include two co-citation hypergraphs: Citeseer (Giles, Bollacker, and Lawrence 1998) and Cora (Sen et al. 2008), and four co-purchasing networks: History, Photo, Computers, and Fitness (Yan et al. 2023). Detailed dataset statistics are presented in Table 1, with comprehensive descriptions provided in Appendix A.1.
Baselines. We compare HiTeC with seven strong SSL baselines covering both graph- and hypergraph-based methods. Since these methods cannot handle raw text directly, we extract features using shallow encoders (SE), PLMs (BERT (Devlin et al. 2019), RoBERTa (Liu et al. 2019)), and graph-agnositc encoder (GIANT (Chien et al. 2022a)). Full details are provided in Appendix A.2.
Evaluation protocol. We evaluate our approach on two hypergraph benchmark tasks: node classification and hyperedge prediction. After pretraining with HiTeC, we follow the linear evaluation protocol (Lee and Shin 2023; Li et al. 2024; Kim et al. 2024a), where the learned node embeddings are fixed and used as inputs to downstream classifiers. Full evaluation details are provided in Appendix A.3.
Implementation details. All baselines use official implementations with default settings. For HiTeC, we tune the contrastive loss weights $\lambda_{e}$ and $\lambda_{s}$ over $\{1,2,3,4\}$. The overlap size $s$ is selected from $\{1,2,3,4,5,10\}$, and the sampling ratio rr% from $\{10\%,20\%,30\%,40\%,50\%\}$. Full implementation details are provided in Appendix A.4.
4.2 Main Results
We evaluate our approach on two hypergraph tasks: node classification and hyperedge prediction (see Appendix B.1). Node classification evaluation. We first evaluate HiTeC on the node classification task. Table 2 presents the results, revealing two key observations: (1) Graph-based SSL methods perform consistently worse than HSSL methods. This is primarily due to the loss of high-order relations caused by clique expansion. Even GIANT, which incorporates textual and structural signals, remains insufficient for modeling high-order structures. These results highlight the necessity of developing SSL methods tailored for TAHG. (2) HiTeC achieves the best overall performance with strong scalability. VilLain performs poorly due to its neglect of textual information, highlighting the critical role of node semantics. TriCL and SE-HSSL also underperform, mainly because of inadequate textual feature extraction and limited modeling of long-range dependencies. While both HyperBERT and HiTeC integrate textual semantics with high-order structure, HyperBERT fails to fully exploit semantic–structural signals and cannot scale to even medium-sized hypergraphs (e.g., History, Photo) due to its joint modeling scheme. In contrast, HiTeC consistently outperforms all baselines, surpassing the second-best method by 9.77% and 10.26% on Cora and Fitness, respectively. This is attributed to the semantic-aware augmentations and $s$-walk-based subgraph contrast, which reduce perturbation noise and enable the model to capture more global dependencies. Furthermore, our two-stage decoupled design brings scalability, with the structure-aware contrastive objective complementing it while preserving high performance.
(a) Impact of components.
(b) Impact of prompt.
Figure 4: Ablation study. (a) Impact of components: “w/o pre” removes the structure-aware pre-training of the text encoder; “w/o shd” replaces the semantic-aware hyperedge drop with the random drop; and “w/o s” removes the subgraph-level contrastive objective. (b) Impact of prompt: “w/o prompt” uses original text only; “w/o domain”, “w/o topology”, and “w/o context” remove the corresponding content from the prompt template.
4.3 Ablation Study
We conduct ablation studies on: (1) key components of HiTeC, (2) prompt template elements, and (3) text encoder backbones (see Appendix B.2).
Impact of components. We first assess the contribution of each component in HiTeC. As shown in Figure 4(a), removing any module results in a performance drop. The most severe drop is caused by “w/o pre”, with 11.25% and 23.41% declines on Computers and Fitness, respectively. This highlights the importance of incorporating topology into the text encoder for learning expressive textual representations. Semantic-aware hyperedge drop is also critical, as its removal (“w/o shd”) leads to a 17.02% drop on Fitness, demonstrating the value of maintaining semantic consistency during structural perturbation.
Impact of prompt. We further analyze the effect of prompt design in Figure 4(b). Overall, our prompt-enhanced augmentation outperforms other variants. “w/o prompt” yields the largest performance drop, confirming the benefits of injecting diverse semantics into text augmentation. Removing individual content, such as context or topology, also leads to noticeable declines, highlighting the importance of incorporating meaningful structural signals into the original text.
4.4 Parameter Sensitivity Analysis
We conduct sensitivity experiments on two key hyperparameters: the overlap size $s$ in the $s$-walk, and the subgraph sampling ratio $r$ (see Appendix B.3).
Figure 5: Sensitive analysis of $s$. The line plot shows model accuracy under different values of $s$. The bar plot indicates the average number of hyperedges in the sampled subgraph, reflecting its exploration range.
Figure 6: Training time vs. accuracy of random walk and $s$-walk. Bars indicate training time, markers denote accuracy.
Overlap size $s$. Figure 5 illustrates how model performance varies with different values of $s$, where model performance first improves and then degrades as $s$ increases. This trend can be attributed to the trade-off between structural coherence and exploration: when the exploration range remains sufficiently broad, a larger $s$ promotes more cohesive subgraphs while preserving essential contextual semantics. However, excessively strict overlap constraints (e.g., ss=10) limit structural diversity and impair performance. Figure 6 further indicates that increasing $s$ leads to smaller subgraphs, thereby accelerating training. These results highlight the importance of selecting an appropriate $s$ to balance accuracy and efficiency.
5 Related Work
Our work is closely related to self-supervised learning on hypergraphs and representation learning on text-attributed graphs (see Appendix C.1).
5.1 Self-supervised Learning on Hypergraphs
SSL provides a promising alternative for hypergraph representation learning without expensive labels. Most existing HSSL methods focus on contrastive learning. HyperGCL (Wei et al. 2022) introduces a learnable augmentation strategy, while TriCL (Lee and Shin 2023) and SE-HSSL (Li et al. 2024) design hierarchical objectives across multiple granularities. Beyond contrastive approaches, generative SSL has also emerged, with HypeBoy (Kim et al. 2024a) proposing a hyperedge-filling task. Other efforts address different settings, such as VilLain (Lee, Lee, and Shin 2024), which targets feature-absent scenarios. While effective in modeling high-order structure, these methods overlook the rich textual semantics in TAHGs. HyperBERT (Bazaga, Lio, and Micklem 2024) pioneers SSL on TAHGs by jointly encoding textual semantics and hypergraph structure, but it incurs high computational overhead and suffers from poor scalability. To bridge the gap, we explore a scalable and effective contrastive learning framework for TAHGs.
6 Conclusion
We propose HiTeC, a hierarchical contrastive learning framework with semantic-aware enhancement, to address limitations of existing HCL methods in TAHGs. To ensure scalability, HiTeC adopts a two-stage decoupled paradigm. The first stage pre-trains a text encoder with a structure-aware contrastive objective, alleviating the graph-agnostic nature of conventional encoders. In the second stage, two semantic-aware augmentation strategies, including prompt-enhanced text augmentation and semantic-aware hyperedge drop, are designed to generate more informative hypergraph views. Furthermore, we introduce a hierarchical contrastive objective, where a novel $s$-walk-based subgraph-level contrast captures long-range structural dependencies beyond node and hyperedge scopes. Extensive experiments on six real-world datasets and multiple downstream tasks demonstrate the effectiveness and scalability of HiTeC.
Appendix
Appendix A Additional Experimental Setup
A.1 Datasets
Since no publicly available benchmark datasets are specifically designed for Text-Attributed Hypergraphs (TAHGs), we construct six datasets spanning two domains: academic co-citation networks and E-commerce platforms. Basic statistics are summarized in Table 1.
Co-citation hypergraphs. We construct two academic hypergraphs, Citeseer (Giles, Bollacker, and Lawrence 1998) and Cora (Sen et al. 2008), based on the standard citation networks, where each node represents a scientific publication. The textual attributes of each node, including the title and abstract, are directly adopted from prior works (He et al. 2024; Wang et al. 2024). To capture high-order relationships, we define a hyperedge as the set of all papers that co-cite the same reference.
E-commerce hypergraphs. We further construct four e-commerce hypergraphs from the Amazon product network (Ni, Li, and McAuley 2019): Book-History, Ele-Photo, Ele-Computers, and Sports-Fitness. In these datasets, each node represents a product, and edges reflect frequent co-purchase or co-browsing behavior. The raw textual attributes of nodes are obtained from the annotations provided by (Yan et al. 2023). Specifically, in Book-History, textual features include book titles and descriptions; in Ele-Computers and Ele-Photo, they consist of high-rated reviews and product summaries; in Sports-Fitness, only product titles are used. To induce high-order structures, we adopt a max-clique-based reconstruction strategy (Bron and Kerbosch 1973; Wang and Kleinberg 2024), where hyperedges are formed by identifying maximal cliques in the co-purchase graph.
A.2 Baselines
We compare our method against seven strong self-supervised learning (SSL) baselines, covering both graph-based and hypergraph-based paradigms.
Graph-based SSL. In this work, we adopt three widely used graph-based SSL frameworks: two contrastive learning methods, GraphCL (You et al. 2020) and BGRL (Thakoor et al. 2022), and one generative method, GraphMAE (Hou et al. 2022). To adapt these methods to hypergraphs, we apply clique expansion following the protocol in (Lee and Shin 2023). For textual feature extraction, we employ shallow encoders (SE) implemented via Skip-Gram (Mikolov et al. 2013), as well as GIANT (Chien et al. 2022a), a graph-agnostic text encoder designed for raw textual attributes.
Hypergraph-based SSL. We include four state-of-the-art hypergraph SSL methods: two contrastive approaches, TriCL (Lee and Shin 2023) and SE-HSSL (Li et al. 2024); one generative method, HypeBoy (Kim et al. 2024a); and VilLain (Lee, Lee, and Shin 2024), which targets the feature-absent setting. All methods adopt the same SE setup and are additionally evaluated with two pre-trained language models: BERT-Base (Devlin et al. 2019) and RoBERTa-Base (Liu et al. 2019).
A.3 Evaluation Protocol
We evaluate our method on two standard hypergraph learning tasks: node classification and hyperedge prediction. After pretraining with HiTeC, we adopt a linear evaluation protocol (Lee and Shin 2023; Li et al. 2024; Kim et al. 2024a), where the pretrained node embeddings are frozen and used as input features for downstream classifiers.
Node classification. Following (Lee and Shin 2023; Li et al. 2024), we randomly split each dataset into 10%/10%/80% for training, validation, and testing. Then, a simple linear classifier is trained on top of the fixed node embeddings using $\ell_{2}$-regularized logistic regression. We report the average test accuracy and standard deviation over 20 random splits, each with 5 different initializations.
Hyperedge prediction. Following (Kim et al. 2024a), we need to generate negative hyperedge samples to evaluate a model on the hyperedge prediction task. In our experiments, we adopt Clique Negative Sampling (CNS) (Yu et al. 2025), a widely used strategy for constructing negative hyperedge samples. Specifically, for a given hyperedge $e$, we replace a node $u\in e$ with a node $v\notin e$ such that $v$ is connected to all other nodes in $e$, i.e., $(e\setminus\{u\})\cup\{v\}$. Then, a two-layer MLP is trained on the fixed node embeddings using the 60%/20%/20% data split. Same as the node classification task, we report the average test accuracy and standard deviation over 20 random splits, each with 5 initializations.
A.4 Implementation Details
All experiments are conducted on a single NVIDIA A5000 GPU (24GB memory). For all baselines, we adopt their official implementations with default settings. For HiTeC, we adopt BERT-Base (Devlin et al. 2019) as the text encoder and apply LoRA (Hu et al. 2022) for parameter-efficient fine-tuning. The text encoder is trained for 5 epochs on small datasets (Citeseer, Cora), 4 epochs on medium datasets (History, Photo), and 3 epochs on large datasets (Computers, Fitness), using a learning rate of 2e-5. The hypergraph encoder is implemented as a single-layer module. We tune the contrastive loss weight $\lambda_{e}$ and $\lambda_{s}$ over $\{1,2,3,4\}$. The overlap threshold $s$ and subgraph sampling ratio $r\%$ are tuned from {1, 2, 3, 4, 5, 10} and {10%, 20%, 30%, 40%, 50%}, respectively. To ensure a fair comparison, we tune the learning rate from {0.01, 0.001, 1e-4, 5e-4} and weight decay from {0, 1e-5} for all methods.
Figure 7: Training time and accuracy under different sampling ratios $r\%$. Bar plots represent training time (in minutes), while line plots indicate the corresponding accuracy. We highlight the relative speedup $\Delta t$ of $r\%{=}30\%$ over $r\%{=}50\%$. All main results are reported with $r\%{=}30\%$.
Method Citeseer Cora History Photo Computers Fitness
Graph SSL
GraphCL + SE 69.48 ± 2.07 59.05 ± 2.16 65.65 ± 1.06 67.48 ± 1.18 68.66 ± 1.33 61.22 ± 1.34
GraphCL + GIANT 72.70 ± 2.06 55.48 ± 1.96 63.02 ± 1.61 68.78 ± 1.16 70.87 ± 0.87 63.02 ± 0.01
BGRL+SE 70.08 ± 2.38 62.45 ± 2.48 64.47 ± 1.18 64.76 ± 1.23 64.66 ± 1.40 60.05 ± 1.05
BGRL + GIANT 72.61 ± 1.99 66.22 ± 2.38 65.13 ± 2.09 65.65 ± 1.10 67.67 ± 0.99 60.92 ± 0.34
GraphMAE + SE 71.71 ± 1.78 64.85 ± 1.66 59.81 ± 0.27 66.33 ± 1.58 64.83 ± 1.72 55.79 ± 0.18
GraphMAE + GIANT 72.25 ± 1.62 66.30 ± 2.19 64.19 ± 0.23 64.83 ± 0.20 66.10 ± 1.20 59.21 ± 0.88
HSSL
TriCL + SE 68.06 ± 1.75 62.43 ± 2.41 67.18 ± 0.24 69.21 ± 0.19 70.63 ± 0.14 63.87 ± 0.30
TriCL + BERT 71.36 ± 1.78 64.02 ± 2.33 70.18 ± 0.31 71.71 ± 0.21 73.20 ± 0.17 65.49 ± 0.40
TriCL + RoBERTa 72.38 ± 1.37 66.49 ± 1.93 68.34 ± 0.58 70.35 ± 0.28 72.20 ± 0.27 63.49 ± 0.17
SE-HSSL + SE 64.80 ± 2.21 65.57 ± 1.88 63.49 ± 1.98 65.42 ± 0.74 66.44 ± 1.09 62.52 ± 0.18
SE-HSSL + BERT 72.29 ± 1.74 65.66 ± 2.25 67.93 ± 0.44 71.27 ± 0.18 70.68 ± 0.91 66.00 ± 0.69
SE-HSSL + RoBERTa 73.09 ± 1.56 65.86 ± 1.94 66.95 ± 0.53 69.65 ± 0.20 71.19 ± 0.16 65.72 ± 0.70
HypeBoy + SE 71.93 ± 2.04 65.17 ± 1.66 OOM OOM OOM OOM
HypeBoy + BERT 74.56 ± 2.10 67.38 ± 1.89 OOM OOM OOM OOM
HypeBoy + RoBERTa 73.47 ± 2.10 66.51 ± 1.87 OOM OOM OOM OOM
VilLain (w/o feat.) 74.25 ± 1.90 63.25 ± 2.57 63.18 ± 0.35 65.02 ± 0.23 65.85 ± 0.28 OOM
HyperBERT 51.84 ± 2.03 45.32 ± 1.47 OOM OOM OOM OOM
HiTeC 74.60 ± 1.84 65.91 ± 2.32 71.83 ± 0.21 72.56 ± 0.17 73.91 ± 0.14 67.32 ± 0.09
Table 3: Evaluation results on hyperedge prediction (% mean ± std). The best and second-best results are highlighted in bold and underline, respectively. ‘w/o feat.” means without node features. OOM denotes out-of-memory on a 24GB GPU.
Dataset DistilBERT BERT RoBERTa DeBERTa Citeseer 67.54 ± 1.49 66.38 ± 1.73 65.09 ± 1.26 64.83 ± 4.65 Cora 74.55 ± 0.02 74.07 ± 0.88 72.32 ± 0.91 70.41 ± 2.24 History 78.05 ± 0.18 79.81 ± 0.16 74.77 ± 0.24 76.20 ± 0.12 Photo 56.90 ± 0.31 59.64 ± 0.16 54.15 ± 0.40 52.54 ± 0.08 Computers 50.64 ± 0.25 55.81 ± 0.12 52.36 ± 0.21 51.48 ± 0.56 Fitness 62.97 ± 0.08 67.18 ± 0.19 64.98 ± 0.78 65.29 ± 0.14
Table 4: Different text encoder backbones.
Appendix B Additional Experiment Results
B.1 Hyperedge Prediction Evaluation
We further evaluate all models on the hyperedge prediction task, with results presented in Table 3. The key observations are as follows: (1) Graph-based SSL methods remain less effective. Their performance is still hindered by the loss of high-order structure due to clique expansion. Although GIANT achieves slight improvements, probably due to its pseudo-labeling strategy for edge-level supervision, it still underperforms compared to HSSL methods enhanced with pre-trained language models. (2) HiTeC achieves the best performance on 5 out of 6 datasets, demonstrating strong generalizability between tasks. TriCL and SE-HSSL underperform due to their limited capacity to capture long-range dependencies. HypeBoy exhibits competitive results on small-scale hypergraphs, benefiting from its hyperedge-filling strategy, but fails to scale to larger datasets, indicating poor scalability. Overall, HiTeC exceeds the second-best method by 1.65% and 1.15% on History and Fitness, respectively, demonstrating its robustness and scalability across diverse hypergraph scenarios.
B.2 Impact of Text Encoder Backbones
We conduct an ablation study on different text encoder backbones, including DistilBERT (Sanh et al. 2019), BERT-Base (Devlin et al. 2019), RoBERTa-Base (Liu et al. 2019), and DeBERTa-Base (He et al. 2020), as shown in Table 4. Among these, BERT achieves the highest overall performance, followed by DistilBERT. These results suggest that selecting a suitable text encoder for each dataset is important. Nevertheless, our framework is compatible with various backbones. In our main experiments, we adopt BERT as the default text encoder backbone.
B.3 Parameter Sensitivity Analysis of Subgraph Sampling Ratio $r\%$
[…]
Appendix C Additional Related Work
C.1 Representation Learning on TAGs
Since hypergraphs can be transformed into graphs via clique expansion, our work is closely related to representation learning on text-attributed graphs (TAGs). Recent TAG methods leverage pre-trained language models to generate numerical features, typically in a cascade (He et al. 2024; Wang et al. 2024) or nested (Yang et al. 2021; Zhao et al. 2023) manner. While most existing TAG methods are supervised, emerging efforts explore self-supervised learning (SSL) to reduce annotation costs. GIANT (Chien et al. 2022a) makes an early attempt by aligning textual and structural features via an extreme multi-label classification objective. More recently, GAugLLM (Fang et al. 2024) proposes an augmentation framework that utilizes large language models (LLMs) to perturb both node features and graph structure, while HASH-CODE (Zhang et al. 2024) formulates a multi-objective contrastive learning strategy tailored for TAGs. However, adapting these methods to hypergraphs via clique expansion incurs substantial structural loss, limiting their effectiveness on TAHGs and motivating the need for dedicated HSSL approaches.

### v3

@@
 Contrastive learning (CL) has become a dominant paradigm for self-supervised hypergraph learning, enabling effective training without costly labels.
-However, node entities in real-world hypergraphs are often associated with rich textual information, which is overlooked in prior works.
-Directly applying existing CL-based methods to such text-attributed hypergraphs (TAHGs) leads to three key limitations: (1) The common use of graph-agnostic text encoders overlooks the correlations between textual content and hypergraph topology, resulting in suboptimal representations.
-(2) Their reliance on random data augmentations introduces noise and weakens the contrastive objective.
-(3) The primary focus on node- and hyperedge-level contrastive signals limits the ability to capture long-range dependencies, which is essential for expressive representation learning.
-Although HyperBERT pioneers CL on TAHGs, its co-training paradigm suffers from poor scalability.
-To fill the research gap, we introduce HiTeC, a two-stage hierarchical contrastive learning framework with semantic-aware augmentation for scalable and effective self-supervised learning on TAHGs.
+However, node entities in real-world hypergraphs are often associated with rich textual information, which has been largely ignored in prior works.
+Directly applying existing CL-based methods to such text-attributed hypergraphs (TAHGs) leads to three key limitations: (1) The common use of graph-agnostic text encoders fails to capture the correlations between textual semantics and hypergraph topology, resulting in less expressive representations.
+(2) Their reliance on random data augmentations introduces noise and weakens the contrastive signals.
+(3) The primary focus on node- and hyperedge-level contrastive signals limits the ability to capture long-range dependencies, which is essential for effective representation learning.
+To address these challenges, we introduce HiTeC, a two-stage hierarchical contrastive learning framework for effective self-supervised learning on TAHGs.
 In the first stage, we pre-train the text encoder with a structure-aware contrastive objective to overcome the graph-agnostic nature of conventional methods.
-In the second stage, we introduce two semantic-aware augmentation strategies, including prompt-enhanced text augmentation and semantic-aware hyperedge drop, to facilitate informative view generation.
-Furthermore, we propose a multi-scale contrastive loss that extends existing objectives with an $s$-walk-based subgraph-level contrast to better capture long-range dependencies.
-By decoupling text encoder pretraining from hypergraph contrastive learning, this two-stage design enhances scalability without compromising representation quality.
+In the second stage, we begin by introducing semantic-aware augmentations, including structure-contextualized text augmentation and semantic-aware hyperedge dropping, to facilitate informative view generation.
+Subsequently, we propose a multi-scale contrastive loss with an $s$-walk-based subgraph-level objective to capture long-range dependencies.
 Extensive experiments on six real-world datasets validate the effectiveness of our proposed method.
+HiTeC: Hierarchical Contrastive Learning on Text-Attributed Hypergraph with Semantic-Aware Augmentation
+Mengting Pan1, Fan Li1, Chen Chen2, Xiaoyang Wang1, Wenjie Zhang1 1The University of New South Wales 2University of Wollongong {mengting.pan,fan.li8,xiaoyang.wang1,wenjie.zhang}@unsw.edu.au chenc@uow.edu.au
 1 Introduction
 Hypergraphs extend traditional graphs by allowing each hyperedge to connect multiple nodes, offering a natural way to model multi-way relationships.
-Such higher-order structures are prevalent in complex real-world systems, e.g., cortical co-activation in brains (Yu et al.
-2011), user-item interactions in e-commerce platforms (Xia, Huang, and Zhang 2022), and multi-author collaborations in academic networks (Kim et al.
-2023).
+Such higher-order structures are prevalent in complex real-world systems, e.g., cortical co-activation in brains Yu et al.
+(2011), user-item interactions in e-commerce platforms Xia et al.
+(2022), and multi-author collaborations in academic networks Kim et al.
+(2023).
 In these scenarios, nodes are often associated with rich textual attributes (e.g., product reviews in user–item networks and paper abstracts in co-authorship networks), which provide essential semantic information for downstream applications.
-While Hypergraph Neural Networks (HNNs) have emerged as a powerful paradigm for hypergraph mining (Chien et al.
-2022b; Kim et al.
-2024b), effectively capturing high-order structural dependencies, they primarily rely on numerical features and task-specific labels.
-This reliance limits their effectiveness in real-world applications, where labeled data is scarce and node attributes are predominantly textual.
-Therefore, underscoring the necessity and significance of self-supervised learning (SSL) on Text-Attributed Hypergraphs (TAHGs).
-Recent advances in hypergraph self-supervised learning (HSSL) have primarily focused on contrastive learning paradigms, which aim to maximize the agreement between two augmented views derived from the original hypergraph.
-HyperGCL (Wei et al.
-2022) pioneers this line of work by employing a learnable augmentation strategy.
-TriCL (Lee and Shin 2023) introduces tri-directional contrastive objectives that model node-, hyperedge-, and membership-level relations in hypergraphs.
-SE-HSSL (Li et al.
-2024) extends this framework by re-designing a hierarchical membership-level objective, achieving state-of-the-art (SOTA) performance.
+While Hypergraph Neural Networks (HNNs) have emerged as a powerful paradigm for hypergraph learning Chien et al.
+(2022b); Kim et al.
+(2024b), effectively capturing high-order structural dependencies, they primarily rely on numerical features and task-specific labels.
+This reliance limits their effectiveness in real-world applications, where labeled data is scarce, and node attributes are predominantly textual.
+This motivates the development of self-supervised learning (SSL) methods tailored to text-attributed hypergraphs (TAHGs) Bazaga et al.
+(2024); Nakajima and Uno (2025).
+Recent advances in hypergraph self-supervised learning (HSSL) primarily focus on the contrastive paradigm, which maximizes agreement between two augmented views derived from the original hypergraph.
+HyperGCL Wei et al.
+(2022) pioneers this direction with a learnable augmentation strategy.
+TriCL Lee and Shin (2023) introduces tri-directional contrastive objectives over node-, hyperedge-, and membership-level relations in hypergraphs.
+SE-HSSL Li et al.
+(2024) further redesigns a hierarchical membership-level objective, achieving state-of-the-art (SOTA) performance.
 Beyond contrastive methods, generative approaches have also gained initial attention.
-HypeBoy (Kim et al.
-2024a) proposes a hyperedge-filling pretext task to learn expressive hypergraph representations.
-Despite recent progress, current HSSL approaches still encounter significant limitations when adopted to hypergraphs with textual attributes: (1) They typically use shallow encoders (e.g., Bag-of-Words (Zhang, Jin, and Zhou 2010) and Skip-Gram (Mikolov et al.
-2013)) or pre-trained language models (e.g., BERT (Devlin et al.
-2019)) to convert raw text into numerical features, ignoring hypergraph structure.
-This graph-agnostic design overlooks the correlation between text and higher-order topology, leading to suboptimal representations (Chien et al.
-2022a).
-(2) Current hypergraph contrastive learning (HCL) methods often apply random augmentations to features and structure, such as randomly masking text or hyperlinks.
-Such perturbations can distort textual meaning and disrupt the alignment between structure and semantic content, thereby introducing noise into the contrastive objective (Fang et al.
-2024).
-(3) Existing HCL methods primarily focus on node- and hyperedge-level signals, while overlooking long-range dependencies and complex topological contexts, which are essential for learning expressive representations (Zhang et al.
-2024).
-HyperBERT (Bazaga, Lio, and Micklem 2024) takes an early attempt at SSL on TAHGs by jointly encoding textual semantics and hypergraph structure via a Transformer-based architecture.
-However, its co-training paradigm incurs substantial memory overhead, limiting scalability even on medium-sized hypergraphs (e.g., History and Photo (Yan et al.
-2023)), as demonstrated in our experiments.
-Moreover, it focuses solely on node-level contrastive signals, ignoring multi-scale structural information that has proven beneficial for HSSL (Lee and Shin 2023; Li et al.
-2024).
-Although recent text-attributed graph studies (Chien et al.
-2022a; Fang et al.
-2024; Zhang et al.
-2024) explore joint modeling of text and structure, they overlook high-order relations in hypergraphs, which may limit their applicability to TAHGs.
-To address these issues, we propose HiTeC, a hierarchical contrastive learning framework with semantic-aware augmentation for scalable and effective SSL on TAHGs.
-HiTeC adopts a two-stage design that decouples the pretraining of the text and hypergraph encoders, significantly improving the scalability over joint training paradigms.
-To overcome the graph-agnostic nature of existing text encoders, we first introduce a structure-aware contrastive objective during text encoder pretraining.
-Positive and negative text pairs are constructed based on structural proximity, encouraging the encoder to capture correlations between textual content and higher-order topology.
-Furthermore, to mitigate the noise introduced by random augmentations, we propose two semantic-aware augmentation strategies: a prompt-enhanced text augmentation module, which injects structural context into raw text to generate diverse yet semantically consistent views, and a semantic-aware hyperedge drop mechanism, which discards hyperedges according to their internal semantic cohesion to preserve meaningful relational patterns.
-Finally, to capture long-range dependencies and multi-scale structural knowledge, we formulate hierarchical contrastive objectives at the node, hyperedge, and subgraph levels.
-Specifically, our subgraph contrastive objective leverages an $s$-walk-based sampling strategy to extract structurally informative subgraphs, facilitating representation learning over broader contexts.
-Our main contributions are summarized as follows:
-  • We propose HiTeC, a scalable and effective two-stage contrastive learning framework for TAHGs, addressing key limitations of existing HSSL methods in modeling complex structural dependencies and rich textual content.
-  • We introduce a structure-aware contrastive objective to pre-train the text encoder, enabling integration of textual content with hypergraph structure, and overcoming the graph-agnostic nature of existing text encoders.
-  • We enhance the hypergraph encoder with semantic-aware augmentations and hierarchical contrastive objectives.
-Specifically, a prompt-enhanced text augmentation and semantic-aware hyperedge drop strategies are designed to reduce perturbation noise.
-And a novel $s$-walk-based subgraph-level signal is employed to capture long-range structural dependencies.
-  • We conduct a comprehensive evaluation of HiTeC on six real-world datasets. The experimental results demonstrate its superiority over strong baselines across various benchmark downstream tasks.
+HypeBoy Kim et al.
+(2024a) proposes a hyperedge-filling pretext task to learn expressive hypergraph representations.
+Despite recent progress, current HSSL approaches still face significant limitations on hypergraphs with textual attributes: (1) They typically use shallow encoders (e.g., Bag-of-Words Zhang et al.
+(2010) and Skip-Gram Mikolov et al.
+(2013)) or pre-trained language models (e.g., BERT Devlin et al.
+(2019)) to convert raw text into numerical features while ignoring hypergraph structure.
+This graph-agnostic design overlooks the correlation between text and higher-order topology, leading to suboptimal representations Chien et al.
+(2022a).
+(2) Current hypergraph contrastive learning (HCL) methods often apply random augmentations to features and structure, e.g., random masking of text or hyperlinks.
+These perturbations may distort textual meaning and disrupt the alignment between structure and semantic content, introducing noise into the contrastive objective Fang et al.
+(2024).
+(3) Existing HCL methods mainly emphasize node- and hyperedge-level signals, but struggle to capture long-range dependencies and richer topological contexts that are crucial for expressive representation learning Zhang et al.
+(2024).
+HyperBERT Bazaga et al.
+(2024) takes an early attempt at SSL on TAHGs by jointly training text and hypergraph encoders through contrastive learning.
+However, its contrastive objective mainly relies on 1-hop hyperedge relations, limiting its ability to capture multi-scale dependencies that have proven beneficial for HSSL Lee and Shin (2023); Li et al.
+(2024).
+Meanwhile, recent text-attributed graph studies Chien et al.
+(2022a); Fang et al.
+(2024); Zhang et al.
+(2024) explore joint text–structure modeling, but their extension to TAHGs remains challenging due to insufficient modeling of higher-order relations inherent to hypergraphs.
+To address these issues, we propose HiTeC, a hierarchical contrastive learning framework for effective SSL on TAHGs.
+HiTeC adopts a two-stage framework that first injects higher-order topology into text representations and then performs semantic-aware hypergraph contrastive learning.
+In the first stage, we pre-train a text encoder with a structure-aware contrastive objective to align textual semantics with higher-order topology, producing topology-enhanced text representations.
+In the second stage, we pre-train a hypergraph encoder with semantic-aware augmentations and hierarchical contrastive objectives.
+Specifically, we introduce structure-contextualized text augmentation and semantic-aware hyperedge dropping to alleviate perturbation noise.
+Based on the augmented views, we further formulate node-, hyperedge-, and subgraph-level contrastive objectives, where an $s$-walk-based subgraph signal captures long-range dependencies and richer structural contexts.
+Our main contributions are as follows:
+  • We propose HiTeC, an effective two-stage contrastive learning framework for TAHGs that addresses key limitations of existing HSSL methods in modeling rich textual attributes and complex hypergraph structure.
+  • We introduce a structure-aware contrastive objective for text encoder pretraining to integrate textual content with hypergraph structure.
+  • We enhance the hypergraph encoder with two semantic-aware augmentation strategies and hierarchical contrastive objectives, where structure-contextualized text augmentation and semantic-aware hyperedge dropping are used to reduce perturbation noise, and an $s$-walk-based subgraph-level contrastive signal captures long-range structural dependencies.
+  • Extensive experiments on six real-world datasets demonstrate the effectiveness and superiority of HiTeC over strong baselines across multiple downstream tasks.
 2 Preliminary
 Notations.
-We define a text-attributed hypergraph (TAHG) as $\mathcal{H}=\{\mathcal{V},\mathcal{E},\mathcal{T}\}$, where $\mathcal{V}=\{v_{i}\}_{i=1}^{|\mathcal{V}|}$ denotes the set of nodes, and $\mathcal{E}=\{e_{j}\}_{j=1}^{|\mathcal{E}|}$ denotes the set of hyperedges, with each hyperedge $e\in\mathcal{E}$ is a non-empty subset of nodes, and $|e|$ is the size of $e$.
-Each node $v_{i}$ is associated with a textual sequence $t_{i}\in\mathcal{T}$, and $\mathcal{T}$ represents the set of all text attributes.
+We define a text-attributed hypergraph (TAHG) as $\mathcal{H}=\{\mathcal{V},\mathcal{E},\mathcal{T}\}$, where $\mathcal{V}=\{v_{i}\}_{i=1}^{|\mathcal{V}|}$ denotes the set of nodes, and $\mathcal{E}=\{e_{j}\}_{j=1}^{|\mathcal{E}|}$ denotes the set of hyperedges, where each hyperedge $e\in\mathcal{E}$ is a non-empty subset of nodes, and $|e|$ is the size of $e$.
+Each node $v_{i}$ is associated with a textual sequence $t_{i}\in\mathcal{T}$, where $\mathcal{T}$ denotes the set of all text attributes.
 Each hyperedge $e_{j}$ is associated with a positive scalar weight $w_{j}$, and all weights formulate a diagonal matrix $\mathbf{W}\in\mathbb{R}^{|\mathcal{E}|\times|\mathcal{E}|}$.
@@
 Hypergraph neural networks.
-Hypergraph neural networks (HNNs) have emerged as effective models for representation learning on hypergraphs (Feng et al.
-2019; Bai, Zhang, and Torr 2021; Chien et al.
-2022b).
+HNNs have emerged as effective tools for representation learning on hypergraphs Feng et al.
+(2019a); Bai et al.
+(2021); Chien et al.
+(2022b).
 Most HNNs adopt a two-stage message passing mechanism that alternates between node-to-hyperedge and hyperedge-to-node aggregation.
@@
 The functions $f^{(l)}_{\mathcal{V}\rightarrow\mathcal{E}}$ and $f^{(l)}_{\mathcal{E}\rightarrow\mathcal{V}}$ are permutation-invariant aggregation functions that propagate messages between nodes and hyperedges.
-This iterative process allows the model to capture complex high-order relations in the hypergraph structure.
+This iterative process allows the model to capture complex higher-order relations in the hypergraph structure.
 Problem statement.
-Given a TAHG $\mathcal{H}=\{\mathcal{V},\mathcal{E},\mathcal{T}\}$, the objective is to learn an optimal function $\Psi:\mathcal{H}\rightarrow(\mathbf{Z}_{\mathcal{V}},\mathbf{Z}_{\mathcal{E}})$ in an unsupervised manner that maps the hypergraph to low-dimensional embeddings of nodes and hyperedges.
+Given a TAHG $\mathcal{H}=\{\mathcal{V},\mathcal{E},\mathcal{T}\}$, the objective is to learn an optimal function $\Psi:\mathcal{H}\rightarrow(\mathbf{Z}_{\mathcal{V}},\mathbf{Z}_{\mathcal{E}})$ in an unsupervised manner, mapping the hypergraph to low-dimensional node and hyperedge embeddings.
 These representations may then be applied to downstream tasks such as node classification and hyperedge prediction.
@@
 Figure 1: The overall framework of HiTeC.
-3.1 Overview
-As illustrated in Figure 1, we present HiTeC, a contrastive learning framework tailored for TAHGs.
-In the first stage, we pre-train the text encoder $f_{\theta}$ with a structure-aware contrastive objective, allowing it to capture topological knowledge and achieve more expressive text representations than graph-agnostic encoders.
-In the second stage, we pre-train a hypegraph encoder $g_{\phi}$ with $f_{\theta}$ frozen.
-To fully exploit the rich semantic context of nodes, we introduce two semantic-aware augmentation strategies on both text- and structure-level, enhancing view diversity while preserving semantic coherence.
-Furthermore, we design a hierarchical contrastive objective across node, hyperedge, and subgraph levels.
-Specifically, the subgraph-level objective, enabled by $s$-walk sampling, captures long-range dependencies.
-By decoupling the training of two encoders, HiTeC achieves better scalability and performance in integrating textual content and high-order structures.
+In this section, we present HiTeC, a novel and effective SSL framework tailored for TAHGs.
+As illustrated in Figure 1, HiTeC adopts a two-stage paradigm.
+In the first stage, a text encoder $f_{\theta}$ is pre-trained with a structure-aware contrastive objective to capture topology-enhanced text representations.
+In the second stage, with $f_{\theta}$ frozen, we pre-train a hypergraph encoder $g_{\phi}$ to fully exploit the rich semantic context of nodes through semantic-aware augmentations and hierarchical contrastive objectives at the node, hyperedge, and subgraph levels.
 3.1 Structure-aware Text Encoder Pretraining
 Text encoder.
-Our text encoder $f_{\theta}$ is built upon a pre-trained language model (PLM) (e.g., BERT (Devlin et al.
-2019)) with a lightweight projection layer.
-To enable efficient training, we apply the parameter-efficient approach, LoRA (Hu et al.
-2022).
-Given a node $v_{i}\in\mathcal{V}$ with textual attribute $t_{i}$, the textual representation is formulated as:
-$x_{i}=f_{\theta}(t_{i}),$ (2)
-where $x_{i}\in\mathbb{R}^{d}$ denotes the embedding corresponding to the [CLS] token in the output sequence.
+Our text encoder $f_{\theta}$ is built upon a pre-trained language model, followed by a two-layer MLP projection head that maps the [CLS] representation to a $d$-dimensional embedding.
+Given a node $v_{i}$ with textual attribute $t_{i}$, the resulting text embedding $x_{i}\in\mathbb{R}^{d}$ is computed as:
+$x_{i}=f_{\theta}(t_{i}).$ (2)
 Structure-aware contrastive objective.
 In real-world hypergraphs, nodes are often semantically closer to their local neighbors than to distant ones.
-For example, in co-citation hypergraphs, papers within the same hyperedge tend to be more topically related.
-To encode these structural priors, we pre-train $f_{\theta}$ by contrasting local and distant semantic contexts.
-Specifically, for each node $v_{i}\in\mathcal{V}$, we treat its 1-hop neighbors $N^{(1)}_{v_{i}}$ as positives, and nodes beyond 1-hop distance $\mathcal{V}\setminus N^{(1)}_{v_{i}}$ as negatives.
-As shown in Figure 1, 1-hop neighbors are defined as nodes belong to the same hyperedges that contain $v_{i}$.
-Contrastive pairs are constructed by mean pooling embeddings from each set:
-$\begin{split}x_{i}^{+}&=\frac{1}{|N^{(1)}_{v_{i}}|}\sum_{v_{j}\in N^{(1)}_{v_{i}}}f_{\theta}(t_{j}),\\ x_{i}^{-}&=\frac{1}{|\mathcal{V}\setminus N^{(1)}_{v_{i}}|}\sum_{v_{j}\in\mathcal{V}\setminus N^{(1)}_{v_{i}}}f_{\theta}(t_{j}),\end{split}$ (3)
-where $x_{i}^{+}\in\mathbb{R}^{d}$ and $x_{i}^{-}\in\mathbb{R}^{d}$ denote the aggregated representations of the positive and negative samples, respectively.
-A triplet margin loss (Schroff, Kalenichenko, and Philbin 2015) is adopted to guide $f_{\theta}$ in distinguishing positive sample $x_{i}^{+}$ from negative sample $x_{i}^{-}$.
-For each anchor $x_{i}$, the loss of the text encoder is defined as:
-$\ell_{t}(x_{i},x_{i}^{+},x_{i}^{-})=\max\{\cos(x_{i},x_{i}^{+})-\cos(x_{i},x_{i}^{-})+m\},$ (4)
-where $\cos(\cdot)$ denotes cosine similarity, and $m$ is the margin hyperparameter. The overall structure-aware objective is averaged across all nodes:
-$\mathcal{L}_{t}=\frac{1}{|\mathcal{V}|}\sum_{v_{i}\in\mathcal{V}}\ell_{t}(x_{i},x_{i}^{+},x_{i}^{-}),$ (5)
-This process allows $f_{\theta}$ to incorporate structural signals into textual representations.
-Figure 2: An example of the prompt template on Cora.
+To encode this structural prior, we treat the 1-hop neighbors $N^{(1)}_{v_{i}}$ of $v_{i}$, i.e., nodes sharing at least one hyperedge with $v_{i}$, as positives, while the remaining nodes $\mathcal{V}\setminus N^{(1)}_{v_{i}}$ are treated as negatives.
+Positive and negative representations are constructed by mean pooling the corresponding text embeddings:
+$\begin{split}x_{i}^{+}&=\frac{1}{|N^{(1)}_{v_{i}}|}\sum_{v_{j}\in N^{(1)}_{v_{i}}}x_{j},\\ x_{i}^{-}&=\frac{1}{|\mathcal{V}\setminus N^{(1)}_{v_{i}}|}\sum_{v_{j}\in\mathcal{V}\setminus N^{(1)}_{v_{i}}}x_{j}.\end{split}$ (3)
+A triplet margin loss Schroff et al.
+(2015) is adopted to guide $f_{\theta}$ in distinguishing positive sample $x_{i}^{+}$ from negative sample $x_{i}^{-}$.
+For each anchor $x_{i}$, the loss is defined as:
+$\ell_{t}(x_{i},x_{i}^{+},x_{i}^{-})=\max\big(\cos(x_{i},x_{i}^{-})-\cos(x_{i},x_{i}^{+})+m,0\big),$ (4)
+where $\cos(\cdot)$ denotes cosine similarity, and $m$ is the margin hyperparameter. The final objective is:
+$\mathcal{L}_{t}=\frac{1}{|\mathcal{V}|}\sum_{v_{i}\in\mathcal{V}}\ell_{t}(x_{i},x_{i}^{+},x_{i}^{-}).$ (5)
 3.2 Semantic-aware Hypergraph Encoder Pretraining
-Existing HCL methods often overlook node textual semantics and rely on random augmentation strategies, resulting in suboptimal performance (Fang et al.
-2024; Wu et al.
-2024).
-Moreover, they struggle to model long-range dependencies, limiting broader contextual understanding (Zhang et al.
-2024; Wu et al.
-2024).
+Existing HCL methods typically overlook node textual semantics and rely on random augmentation strategies, resulting in suboptimal performance Fang et al.
+(2024); Wu et al.
+(2024).
+Moreover, they struggle to model long-range dependencies, limiting broader contextual understanding Zhang et al.
+(2024); Wu et al.
+(2024).
 To tackle these issues, we propose a semantic-aware contrastive learning strategy to pre-train $g_{\phi}$.
@@
 Semantic-aware augmentation.
-To mitigate the noise from random perturbations, we introduce two novel augmentation techniques: prompt-enhanced text augmentation, which generates diverse yet semantically consistent textual views, and semantic-aware hyperedge drop, which ensures structural perturbations remain aligned with the semantics.
-Prompt-enhanced text augmentation.
-As feature-level perturbations can easily distort semantic meaning, we instead focus on augmenting the raw text using a prompt-enhanced strategy.
-For each node text $t_{i}$, we construct an augmented view $\tilde{t}_{1,i}=\text{Prompt}(t_{i},\mathcal{H})$, while retaining the original text as $\tilde{t}_{2,i}=t_{i}$.
-As shown in Figure 2,the prompt template integrates the original text with domain knowledge, global topology, and neighbor context, concatenated into a single sequence.
-Each component adds a distinct perspective: domain knowledge offers semantic background, global topology introduces structural awareness, and neighbor context provides complementary local semantics.
-This simple yet effective strategy generates semantically consistent but contextually diverse text variants, with a shared $f_{\theta}$ that maps them to a unified space, denoted as $\tilde{\mathbf{X}}_{1}$ and $\tilde{\mathbf{X}}_{2}$, respectively.
-Semantic-aware hyperedge drop.
-Effective structure augmentation should consider both textual semantics and the original topology.
-To achieve that, we define a semantic cohesiveness score that reflects both the structural grouping of nodes and their semantic similarity.
-This score is then used to guide the drop probability of each node–hyperedge incidence: connections within more cohesive hyperedges are more likely to be preserved.
-In this way, our strategy encourages the retention of structurally and semantically aligned patterns, while selectively pruning noisy or weakly related hyperlinks.
-We formally define the it as follows.
+To eliminate the impact of noise arising from random perturbation, we introduce two semantic-aware augmentation strategies: structure-contextualized text augmentation and semantic-aware hyperedge dropping.
+Structure-contextualized text augmentation.
+As perturbations applied directly to text embeddings can easily distort semantic meaning, we instead augment raw text with structure-aware contextual information.
+Specifically, for each node text attribute $t_{i}$, we first construct a structure-contextualized text view by concatenating three contextual components with the original text: $\tilde{t}_{1,i}=[\texttt{[DK]}\,||\,\texttt{[GT]}\,||\,\texttt{[NC]}\,||\,t_{i}]$, where: (i) [DK] denotes domain knowledge (e.g., academic or e-commerce) shared across all nodes; (ii) [GT] encodes global topological information for node $v_{i}$ (e.g., node degree and local sparsity); and (iii) [NC] represents neighbor context, constructed by concatenating the raw texts of sampled neighbors of $v_{i}$.
+The second text view is the original text, i.e., $\tilde{t}_{2,i}=t_{i}$.
+This simple yet effective strategy generates semantically consistent but contextually diverse text variants, which are encoded by the shared encoder $f_{\theta}$ into a unified embedding space, denoted as $\tilde{\mathbf{X}}_{1}$ and $\tilde{\mathbf{X}}_{2}$, respectively.
+The full template for constructing the structure-contextualized text view is provided in Appendix A.
+Semantic-aware hyperedge dropping.
+Effective structure augmentation should consider both textual semantics and topology.
+To achieve this, we define a semantic cohesiveness score that reflects both the structural grouping of nodes and their semantic similarity, which is further used to guide the drop probability of each node–hyperedge incidence.
+Connections within more cohesive hyperedges are more likely to be preserved, encouraging the retention of structurally and semantically aligned patterns, while selectively pruning noisy or weakly node–hyperedge connections.
 Definition 1 (semantic cohesiveness score).
-Given a TAHG $\mathcal{H}(\mathcal{V},\mathcal{E},\mathbf{X})$, the semantic cohesiveness score of a hyperedge $e\in\mathcal{E}$ is defined as the average pairwise cosine similarity among its member nodes.
+Given a TAHG $\mathcal{H}(\mathcal{V},\mathcal{E},\mathbf{X})$, the semantic cohesiveness score of a hyperedge $e\in\mathcal{E}$ is defined as the average pairwise similarity among its member nodes.
 Let $\mathbf{X}_{e}\in\mathbb{R}^{|e|\times d}$ denote the normalized feature matrix for nodes in $e$.
@@
 $s(e)=\frac{1}{\binom{|e|}{2}}\sum_{i=1}^{|e|}\sum_{j=i+1}^{|e|}\bm{S}_{e}(i,j),\quad\bm{S}_{e}=\mathbf{X}_{e}\mathbf{X}_{e}^{\top},$
-where $\bm{S}_{e}(i,j)$ denotes the similarity between the $i$-th and $j$-th node in hyperedge $e$.
-The drop probability of each node-hyperedge incidence $(v_{i},e_{j})$ where $v_{i}\in e_{j}$ is then guided by $s(e_{j})$:
+where $\bm{S}_{e}(i,j)$ denotes the similarity between the $i$-th and $j$-th nodes in hyperedge $e$.
+The drop probability of each incidence $(v_{i},e_{j})$, where $v_{i}\in e_{j}$, is then determined by $s(e_{j})$:
 $p_{\text{drop}}(v_{i},e_{j})=1-\sigma\left(\frac{s(e_{j})-0.5}{\tau_{\text{drop}}}\right),$ (6)
 where $\sigma(\cdot)$ is the sigmoid function and $\tau_{\text{drop}}$ is a temperature hyperparameter.
-Higher cohesiveness scores lead to lower drop probabilities.
-In implementation, we construct a binary masking matrix $\mathbf{M}\in\{0,1\}^{|\mathcal{V}|\times|\mathcal{E}|}$, where each entry $m_{ij}\sim\mathcal{B}(1-p_{\text{drop}}(v_{i},e_{j}))$ is sampled from a Bernoulli distribution.
-The perturbed incidence matrix $\tilde{\mathbf{H}}$ is represented as:
+Higher cohesiveness scores correspond to lower drop probabilities.
+In implementation, we construct a binary masking matrix $\mathbf{M}\in\{0,1\}^{|\mathcal{V}|\times|\mathcal{E}|}$, where entries $m_{ij}\sim\mathcal{B}(1-p_{\text{drop}}(v_{i},e_{j}))$ are sampled from a Bernoulli distribution.
+The perturbed matrix $\tilde{\mathbf{H}}$ is:
 $\tilde{\mathbf{H}}=\mathbf{M}\odot\mathbf{H},$ (7)
@@
 Hypergraph encoder.
-We take the HGNN (Feng et al.
-2019) with the element-wise mean pooling layer as our hypergraph encoder.
+We take the HGNN Feng et al.
+(2019b) with the element-wise mean pooling layer as our hypergraph encoder.
 At the $l$-th layer, the hyperedge representation $\mathbf{Z}_{\mathcal{E}}^{(l)}\in\mathbb{R}^{|\mathcal{E}|\times d}$ and the node representation $\mathbf{Z}_{\mathcal{V}}^{(l)}\in\mathbb{R}^{|\mathcal{V}|\times d}$ are updated as:
 $\begin{split}\mathbf{Z}_{\mathcal{E}}^{(l)}&=\rho\left(\mathbf{D}_{e}^{-1}\mathbf{H}^{\top}\mathbf{Z}_{\mathcal{V}}^{(l-1)}\mathbf{\Theta}_{\mathcal{E}}^{(l)}+\bm{b}_{e}^{(l)}\right),\\ \mathbf{Z}_{\mathcal{V}}^{(l)}&=\rho\left(\mathbf{D}_{v}^{-1}\mathbf{H}\mathbf{W}\mathbf{Z}_{\mathcal{E}}^{(l)}\mathbf{\Theta}_{\mathcal{V}}^{(l)}+\bm{b}_{v}^{(l)}\right),\end{split}$ (8)
-where $\rho(\cdot)$ is the PReLU activation function, $\mathbf{\Theta}_{\mathcal{E}}^{(l)}$ and $\mathbf{\Theta}_{\mathcal{V}}^{(l)}$ are trainable weight matrices, and $\bm{b}_{e}^{(l)}$ and $\bm{b}_{v}^{(l)}$ are corresponding bias terms.
-$\mathbf{W}$ is initialized as an identity matrix, which means equal weights for all hyperedges.
+where $\rho(\cdot)$ is the PReLU activation function, $\mathbf{\Theta}_{\mathcal{E}}^{(l)}$ and $\mathbf{\Theta}_{\mathcal{V}}^{(l)}$ are trainable weight matrices, and $\bm{b}_{e}^{(l)}$ and $\bm{b}_{v}^{(l)}$ are bias terms.
+$\mathbf{W}$ is initialized as an identity matrix, assigning equal weights to all hyperedges.
 Hierarchical contrastive objectives.
-Prior works (Lee and Shin 2023; Li et al.
-2024) have demonstrated that multi-scale contrastive objectives are beneficial for HSSL.
-To better capture structural semantics at different granularities, we adopt a hierarchical contrastive learning objective spanning node, hyperedge, and subgraph levels.
-In particular, the novel subgraph-level objective, based on $s$-walk sampling, enables the model to capture long-range dependencies beyond local structure.
-All objectives are optimized with InfoNCE loss (Oord, Li, and Vinyals 2018).
+Prior works Lee and Shin (2023); Li et al.
+(2024) have shown that multi-scale contrastive objectives are beneficial for HSSL.
+To better capture structural semantics at different granularities, we adopt a hierarchical contrastive learning framework across node, hyperedge, and subgraph levels.
+In particular, we introduce a novel subgraph-level objective based on $s$-walk sampling to capture long-range dependencies beyond local structures.
+All objectives are optimized using the widely adopted contrastive loss, InfoNCE Oord et al.
+(2018).
 […]
 $\ell_{n}(\mathbf{z}_{1,v_{i}},\mathbf{z}_{2,v_{i}})=-\log\frac{e^{\cos(\mathbf{z}_{1,v_{i}},\mathbf{z}_{2,v_{i}})/\tau_{n}}}{\sum_{k=1}^{|\mathcal{V}|}e^{\cos(\mathbf{z}_{1,v_{i}},\mathbf{z}_{2,v_{k}})/\tau_{n}}},$ (9)
-where $\tau_{n}$ is the temperature coefficient. In practice, we symmetrize the loss by using the second view’s embedding as the anchor and obtain the final node-level contrastive loss:
+where $\tau_{n}$ is the temperature coefficient.
+In practice, we further symmetrize the contrastive loss by using the second view’s embedding as the anchor, yielding the final node-level contrastive objective:
 $\mathcal{L}_{n}=\frac{1}{2|\mathcal{V}|}\sum_{i=1}^{|\mathcal{V}|}\left[\ell_{n}(\mathbf{z}_{1,v_{i}},\mathbf{z}_{2,v_{i}})+\ell_{n}(\mathbf{z}_{2,v_{i}},\mathbf{z}_{1,v_{i}})\right].$ (10)
@@
 $\ell_{\text{e}}(\mathbf{z}_{1,e_{j}},\mathbf{z}_{2,e_{j}})=-\log\frac{e^{\cos(\mathbf{z}_{1,e_{j}},\mathbf{z}_{2,e_{j}})/\tau_{e}}}{\sum_{k=1}^{|\mathcal{E}|}e^{\cos(\mathbf{z}_{1,e_{j}},\mathbf{z}_{2,e_{k}})/\tau_{e}}},$ (11)
-where $\tau_{e}$ is the temperature coefficient. Symmetrizing the loss, the final hyperedge-level objective can be written as:
+where $\tau_{e}$ is the temperature coefficient. By symmetrizing the contrastive loss, the final hyperedge-level objective can be written as:
 $\mathcal{L}_{e}=\frac{1}{2|\mathcal{E}|}\sum_{j=1}^{|\mathcal{E}|}\left[\ell_{e}(\mathbf{z}_{1,e_{j}},\mathbf{z}_{2,e_{j}})+\ell_{e}(\mathbf{z}_{2,e_{j}},\mathbf{z}_{1,e_{j}})\right].$ (12)
@@
 Figure 2: An example of the $s$-walk-based subgraph sampling process.
-For center node $v_{c}$, we perform $s$-walk with $s=2$ and $l=4$ on $\mathcal{H}$.
-The hyperedge sequences visited are $e_{1},e_{3},e_{4},e_{5}]$.
-The sampled subgraph is denoted as $\mathcal{H}_{s_{c}}$.
-While node- and hyperedge-level objectives capture local semantic and structural signals, they may fail to model long-range dependencies in hypergraphs.
-To bridge this gap, we further introduce a subgraph-level objective by constructing broader structural signals through a novel $s$-walk-based sampling strategy.
-Unlike random walk, $s$-walk is a high-order sequence in hypergraphs, where $s$ measures the overlap strength between hyperedges, enabling the extraction of cohesive and semantically meaningful substructures unique to hypergraphs (Aksoy et al.
-2020; Preti, De Francisci Morales, and Bonchi 2024).
-We next define $s$-walk and detail our subgraph sampling method.
+Given the center node $v_{c}$, we perform an $s$-walk with $s=2$ and $l=3$ on $\mathcal{H}$.
+The resulting hyperedge traversal sequence is $e_{1},e_{3},e_{4}]$, and the sampled subgraph is denoted as $\mathcal{H}_{s_{c}}$.
+While node- and hyperedge-level objectives effectively capture local structural signals, they remain insufficient for modeling long-range dependencies in hypergraphs.
+We therefore introduce a subgraph-level objective based on an $s$-walk sampling strategy.
+Unlike random walks, an $s$-walk defines a high-order traversal sequence in hypergraphs, where $s$ controls the overlap strength between consecutive hyperedges, enabling the extraction of cohesive and semantically meaningful substructures unique to hypergraphs Aksoy et al.
+(2020); Preti et al.
+(2024).
+We next formally define $s$-walk and describe our subgraph sampling method in detail.
 Definition 2 ($s$-walk).
-Given $s\in\mathbb{N}^{+}$, an $s$-walk of length $l$ is a sequence of hyperedges $e_{1},\ldots,e_{l}]$ such that each consecutive pair of hyperedges shares at least $s$ nodes, i.e., $|e_{j}\cap e_{j+1}|\geq s$ for all $j\in[1,l-1]$.
-As illustrated in Figure 3, we construct an $s$-walk-based subgraph $\mathcal{H}_{s_{c}}=\{\mathcal{V}_{s_{c}},\mathcal{E}_{s_{c}}\}$ by first selecting a starting hyperedge $e_{1}$ that contains the center node $v_{c}$, and then performing an $s$-walk of length $l$.
-We randomly sample one node from each visited hyperedge.
-The sampled subgraph includes all visited hyperedges and their associated nodes: $\mathcal{E}_{s_{c}}=\{e_{j}\}_{j=1}^{l}$, $\mathcal{V}_{s_{c}}=\bigcup_{j=1}^{l}e_{j}$.
-Each subgraph $\mathcal{H}_{s_{c}}$ can be encoded by the shared hypergraph encoder $g_{\phi}(\cdot)$, and its representation is obtained via mean pooling $P_{\text{mean}}$ over the node embeddings:
+Given $s\in\mathbb{N}^{+}$, an $s$-walk of length $l$ is defined as an ordered sequence of hyperedges $e_{1},\ldots,e_{l}]$, where every consecutive pair of hyperedges shares at least $s$ common nodes, i.e., $|e_{j}\cap e_{j+1}|\geq s$ for all $j\in[1,l-1]$.
+As illustrated in Figure 2, an $s$-walk-based subgraph $\mathcal{H}_{s_{c}}=(\mathcal{V}_{s_{c}},\mathcal{E}_{s_{c}})$ is constructed from a center node $v_{c}$ via iterative $s$-walk sampling.
+We first randomly select a hyperedge $e_{1}$ containing $v_{c}$.
+At each step, we identify hyperedges sharing at least $s$ nodes with the current hyperedge $e_{j}$, randomly sample one as $e_{j+1}$, and then randomly choose a node from $e_{j+1}$ as the transition node.
+The process stops when the traversal length reaches $l$.
+The sampled subgraph consists of the visited hyperedges $\mathcal{E}_{s_{c}}=\{e_{j}\}_{j=1}^{l}$ and their associated nodes $\mathcal{V}_{s_{c}}=\bigcup_{j=1}^{l}e_{j}$.
+Each sampled subgraph $\mathcal{H}_{s_{c}}$ is encoded by the shared hypergraph encoder $g_{\phi}(\cdot)$, and its representation is obtained through mean pooling $P_{\text{mean}}$ over the corresponding node embeddings:
 $\mathbf{z}_{s_{c}}=P_{\text{mean}}\!\left(g_{\phi}(\mathcal{H}_{s_{c}})\right).$ (13)
 However, constructing and encoding subgraphs for all nodes incurs a high computational cost.
-To mitigate this, we follow prior findings suggesting that high-degree nodes tend to carry richer contextual information (Fang et al.
-2025).
-Specifically, we select a subset of representative nodes $\mathcal{V}_{r}\subseteq\mathcal{V}$ by ranking node degrees and retaining the top $r\%$.
+To mitigate this, we follow prior findings suggesting that high-degree nodes tend to carry richer contextual information Fang et al.
+(2025).
+Specifically, we select a representative node subset $\mathcal{V}_{r}\subseteq\mathcal{V}$ by ranking node degrees and retaining the top $r\%$.
 As validated in our experiments, this strategy maintains competitive performance even with a significantly reduced number of subgraphs.
@@
 $\mathcal{L}_{\text{s}}=\frac{1}{2|\mathcal{V}_{r}|}\sum_{m=1}^{|\mathcal{V}_{r}|}\left[\ell_{s}(\mathbf{z}_{1,s_{m}},\mathbf{z}_{2,s_{m}})+\ell_{\text{s}}(\mathbf{z}_{2,s_{m}},\mathbf{z}_{1,s_{m}})\right].$ (15)
-Overall objective. Finally, by integrating Eq. (10), (12), and (15), our proposed contrastive loss is formulated as:
+Overall objective. Combining Eq. (10), (12), and (15), the overall contrastive loss is defined as: contrastive loss is defined as:
 $\mathcal{L}=\mathcal{L}_{n}+\lambda_{e}\mathcal{L}_{e}+\lambda_{s}\mathcal{L}_{s},$ (16)
-where $\lambda_{e}$ and $\lambda_{s}$ are balancing weights.
-This hierarchical contrastive framework enables the model to capture patterns at multiple granularities, generating more robust and discriminative node representations.
+where $\lambda_{e}$ and $\lambda_{s}$ are balancing weights. A detailed complexity analysis is provided in Appendix B.
 4 Experiments
@@
 Method Citeseer Cora History Photo Computers Fitness
-Graph SSL
-GraphCL + SE 59.63 ± 1.58 56.69 ± 1.99 62.86 ± 0.22 42.61 ± 0.21 34.45 ± 0.38 52.11 ± 0.43
-GraphCL + GIANT 57.58 ± 1.90 55.64 ± 1.50 62.08 ± 0.24 42.74 ± 0.22 34.65 + 0.23 52.04 ± 0.34
+Graph SSL GraphCL + SE 59.63 ± 1.58 56.69 ± 1.99 62.86 ± 0.22 42.61 ± 0.21 34.45 ± 0.38 52.11 ± 0.43
+GraphCL + GIANT 57.58 ± 1.90 55.64 ± 1.50 62.08 ± 0.24 42.74 ± 0.22 34.65 ± 0.23 52.04 ± 0.34
 BGRL + SE 60.81 ± 1.45 59.72 ± 1.32 65.74 ± 0.30 44.60 ± 0.30 39.40 ± 0.46 55.66 ± 0.33
@@
 GraphMAE + GIANT 50.11 ± 1.74 53.59 ± 1.07 56.20 ± 0.12 41.54 ± 0.06 25.57 ± 0.12 42.04 ± 0.04
-HSSL
-TriCL + SE 60.78 ± 1.00 58.79 ± 1.30 69.28 ± 0.32 47.11 ± 0.44 43.43 ± 0.66 45.46 ± 0.21
+HSSL TriCL + SE 60.78 ± 1.00 58.79 ± 1.30 69.28 ± 0.32 47.11 ± 0.44 43.43 ± 0.66 45.46 ± 0.21
 TriCL + BERT 64.50 ± 1.08 62.67 ± 1.13 75.01 ± 0.22 51.25 ± 0.62 46.60 ± 0.90 48.97 ± 0.63
@@
 HypeBoy + RoBERTa 50.41 ± 2.88 30.84 ± 0.95 OOM OOM OOM OOM
-VilLain (w/o feat.) 42.22 ± 1.72 53.72 ± 1.00 40.99 ± 3.06 33.94 ± 1.39 20.99 ± 0.91 OOM
-HyperBERT 30.25 ± 3.04 26.85 ± 0.66 OOM OOM OOM OOM
-HiTeC 66.38 ± 1.73 74.07 ± 0.88 79.81 ± 0.16 59.64 ± 0.16 55.81 ± 0.12 67.18 ± 0.19
-Table 2: Node classification accuracy (% mean ± std) across six datasets.
+VilLain (w/o feat.) 42.22 ± 1.72 53.72 ± 1.00 40.99 ± 3.06 33.94 ± 1.39 20.99 ± 0.91 42.04 ± 0.04
+HyperBERT 58.25 ± 3.04 47.03 ± 0.77 62.12 ± 0.19 41.46 ± 0.14 26.79 ± 0.20 OOM
+HiTeC 68.69 ± 0.97 74.07 ± 0.88 79.81 ± 0.16 59.64 ± 0.16 55.81 ± 0.12 67.18 ± 0.19
+Table 2: Evaluation results on node classification task(% mean $\pm$ std).
 The best and second-best results are highlighted in bold and underline, respectively.
-“w/o feat.” means without node features.
-OOM denotes out-of-memory on a 24GB GPU.
+“w/o feat.” denotes methods without node features.
+OOM indicates that the method runs out of memory on a single 80GB GPU.
 4.1 Experimental Setup
 Datasets.
-Since there are no publicly available TAHG datasets, we introduce six new text-attributed hypergraph datasets using a max clique-based hypergraph reconstruction method (Bron and Kerbosch 1973; Wang and Kleinberg 2024).
-These include two co-citation hypergraphs: Citeseer (Giles, Bollacker, and Lawrence 1998) and Cora (Sen et al.
-2008), and four co-purchasing networks: History, Photo, Computers, and Fitness (Yan et al.
-2023).
-Detailed dataset statistics are presented in Table 1, with comprehensive descriptions provided in Appendix A.1.
+As no public TAHG datasets are available, we construct six text-attributed hypergraph datasets via max clique-based hypergraph reconstruction Bron and Kerbosch (1973); Wang and Kleinberg (2024).
+They include two co-citation hypergraphs (Citeseer Giles et al.
+(1998) and Cora Sen et al.
+(2008)) and four co-purchasing networks (History, Photo, Computers, and Fitness Yan et al.
+(2023)).
+Dataset statistics are summarized in Table 1, with details in Appendix C.1.
 Baselines.
-We compare HiTeC with seven strong SSL baselines covering both graph- and hypergraph-based methods.
-Since these methods cannot handle raw text directly, we extract features using shallow encoders (SE), PLMs (BERT (Devlin et al.
-2019), RoBERTa (Liu et al.
-2019)), and graph-agnositc encoder (GIANT (Chien et al.
-2022a)).
-Full details are provided in Appendix A.2.
+We compare HiTeC with eight representative SSL baselines: (1) three graph-based methods (GraphCL You et al.
+(2020), BGRL Thakoor et al.
+(2022), and GraphMAE Hou et al.
+(2022)); (2) four hypergraph-based methods (TriCL Lee and Shin (2023), SE-HSSL Li et al.
+(2024), HypeBoy Kim et al.
+(2024a), and VilLain Lee et al.
+(2024)); and (3) one TAHG-based method, HyperBERT Bazaga et al.
+(2024).
+Since graph- and hypergraph-based methods cannot process raw text directly, we extract node features using different text encoders.
+Specifically, both graph- and hypergraph-based methods use shallow encoders (SE), i.e., Skip-Gram Mikolov et al.
+(2013).
+Graph-based methods additionally use GIANT Chien et al.
+(2022a), a self-supervised encoder for text-attributed graphs (TAGs), while hypergraph-based methods use pretrained language models, including BERT Devlin et al.
+(2019) and RoBERTa Liu et al.
+(2019).
+Detailed descriptions of the baselines are provided in Appendix C.2.
 Evaluation protocol.
 We evaluate our approach on two hypergraph benchmark tasks: node classification and hyperedge prediction.
-After pretraining with HiTeC, we follow the linear evaluation protocol (Lee and Shin 2023; Li et al.
-2024; Kim et al.
-2024a), where the learned node embeddings are fixed and used as inputs to downstream classifiers.
-Full evaluation details are provided in Appendix A.3.
+After pretraining with HiTeC, we adopt the linear evaluation protocol Lee and Shin (2023); Li et al.
+(2024); Kim et al.
+(2024a), where the learned node embeddings are frozen and used as input features for downstream classifiers.
+For node classification, we train a logistic regression classifier with a 10%/10%/80% train/validation/test split.
+For hyperedge prediction, following Kim et al.
+(2024a), we sample an equal number of negative hyperedges to match the ground-truth positives and split the data using a 60%/20%/20% ratio, after which a two-layer MLP classifier is trained for prediction.
+For each dataset, we report the average test accuracy and standard deviation over 20 random data splits, each evaluated with 5 different random initializations.
+Full evaluation details are provided in Appendix C.3.
 Implementation details.
-All baselines use official implementations with default settings.
-For HiTeC, we tune the contrastive loss weights $\lambda_{e}$ and $\lambda_{s}$ over $\{1,2,3,4\}$.
-The overlap size $s$ is selected from $\{1,2,3,4,5,10\}$, and the sampling ratio rr% from $\{10\%,20\%,30\%,40\%,50\%\}$.
-Full implementation details are provided in Appendix A.4.
+For HiTeC, we adopt BERT-Base Devlin et al.
+(2019) as the default text encoder and apply LoRA Hu et al.
+(2022) for parameter-efficient fine-tuning.
+We further evaluate other text encoder backbones in Appendix D.2.
+We tune the $s$-walk length $l$ and overlap size $s$ in $\{1,2,3,4,5,10\}$, the sampling ratio $r$ in $\{10,20,30,40,50\}$, and $\lambda_{e}$ and $\lambda_{s}$ in $\{1,2,3,4\}$.
+Additional implementation details are shown in Appendix C.4.
 4.2 Main Results
-We evaluate our approach on two hypergraph tasks: node classification and hyperedge prediction (see Appendix B.1).
+In this section, we evaluate our approach on two representative hypergraph tasks: node classification and hyperedge prediction (see Appendix D.1).
 Node classification evaluation.
-We first evaluate HiTeC on the node classification task.
-Table 2 presents the results, revealing two key observations: (1) Graph-based SSL methods perform consistently worse than HSSL methods.
-This is primarily due to the loss of high-order relations caused by clique expansion.
-Even GIANT, which incorporates textual and structural signals, remains insufficient for modeling high-order structures.
-These results highlight the necessity of developing SSL methods tailored for TAHG.
-(2) HiTeC achieves the best overall performance with strong scalability.
-VilLain performs poorly due to its neglect of textual information, highlighting the critical role of node semantics.
-TriCL and SE-HSSL also underperform, mainly because of inadequate textual feature extraction and limited modeling of long-range dependencies.
-While both HyperBERT and HiTeC integrate textual semantics with high-order structure, HyperBERT fails to fully exploit semantic–structural signals and cannot scale to even medium-sized hypergraphs (e.g., History, Photo) due to its joint modeling scheme.
+Table 2 presents the node classification results, revealing two key observations: (1) Graph-based SSL methods consistently underperform HSSL methods due to the loss of high-order relations caused by clique expansion.
+Even the TAG encoder GIANT, which jointly encodes text and pairwise structure, remains insufficient for modeling higher-order structures, highlighting the need for dedicated SSL methods for TAHGs.
+(2) HiTeC achieves the best overall performance across all datasets.
+Methods that neglect rich node features (e.g., VilLain) achieve limited performance, highlighting the importance of textual semantics in TAHG learning.
+HCL methods (e.g., TriCL and SE-HSSL) rely on structure-driven contrastive signals without semantic-aware augmentation, resulting in suboptimal performance.
+While HyperBERT jointly models textual and hypergraph information, it lacks multi-scale structural and semantic contrastive signals.
 In contrast, HiTeC consistently outperforms all baselines, surpassing the second-best method by 9.77% and 10.26% on Cora and Fitness, respectively.
-This is attributed to the semantic-aware augmentations and $s$-walk-based subgraph contrast, which reduce perturbation noise and enable the model to capture more global dependencies.
-Furthermore, our two-stage decoupled design brings scalability, with the structure-aware contrastive objective complementing it while preserving high performance.
-(a) Impact of components.
-(b) Impact of prompt.
+These improvements are mainly attributed to semantic-aware augmentations, which reduce perturbation noise, and the $s$-walk-based subgraph signals, which capture more informative high-order dependencies.
+Additional analysis on efficiency is provided in Appendix D.5.
+4.3 Ablation Study
+(a) Key component ablation.
+(b) Text view ablation.
 Figure 3: Ablation study.
-(a) Impact of components: “w/o pre” removes the structure-aware pre-training of the text encoder; “w/o shd” replaces the semantic-aware hyperedge drop with the random drop; and “w/o s” removes the subgraph-level contrastive objective.
-(b) Impact of prompt: “w/o prompt” uses original text only; “w/o domain”, “w/o topology”, and “w/o context” remove the corresponding content from the prompt template.
-4.3 Ablation Study
-We conduct ablation studies on: (1) key components of HiTeC, (2) prompt template elements, and (3) text encoder backbones (see Appendix B.2).
-Impact of components.
-We first assess the contribution of each component in HiTeC.
-As shown in Figure 4(a), removing any module results in a performance drop.
-The most severe drop is caused by “w/o pre”, with 11.25% and 23.41% declines on Computers and Fitness, respectively.
-This highlights the importance of incorporating topology into the text encoder for learning expressive textual representations.
-Semantic-aware hyperedge drop is also critical, as its removal (“w/o shd”) leads to a 17.02% drop on Fitness, demonstrating the value of maintaining semantic consistency during structural perturbation.
-Impact of prompt.
-We further analyze the effect of prompt design in Figure 4(b).
-Overall, our prompt-enhanced augmentation outperforms other variants.
-“w/o prompt” yields the largest performance drop, confirming the benefits of injecting diverse semantics into text augmentation.
-Removing individual content, such as context or topology, also leads to noticeable declines, highlighting the importance of incorporating meaningful structural signals into the original text.
+(a) Key component ablation: “w/o pre” removes structure-aware text encoder pre-training; “w/o shd” replaces semantic-aware hyperedge dropping with random drop; “w/o sct” removes the structure-contextualized text view; and “w/o s” removes the subgraph-level objective.
+(b) Text view ablation: “w/o DK”, “w/o GT”, and “w/o NC” remove domain knowledge, global topology, and neighbor context from the structure-contextualized text view, respectively.
+In this section, we conduct ablation studies on: (1) key components of HiTeC, (2) structure-contextualized text view components.
+Impact of key components.
+We evaluate the contribution of key components in HiTeC.
+As shown in Figure 3(a), removing any module noticeably degrades performance.
+The most significant drop is observed when removing text encoder pre-training (‘w/o pre”), resulting in a decrease of 13.41% on Fitness, highlighting the effectiveness of injecting topological information into the text encoder.
+The subgraph-level objective also plays a critical role: removing it (‘w/o s”) causes a decrease of 7.14% on Photo, demonstrating the benefit of modeling long-range dependencies inherent in TAHGs.
+Impact of structure-contextualized text view components.
+We further evaluate the effectiveness of different components in the structure-contextualized text view in Figure 3(b).
+Overall, the complete view consistently outperforms all alternative variants.
+Specifically, removing neighbor context (“w/o NC”) and global topology (“w/o GT”) leads to pronounced performance degradation, underscoring the importance of incorporating meaningful structural signals into the original text.
 4.4 Parameter Sensitivity Analysis
-We conduct sensitivity experiments on two key hyperparameters: the overlap size $s$ in the $s$-walk, and the subgraph sampling ratio $r$ (see Appendix B.3).
-Figure 5: Sensitive analysis of $s$.
-The line plot shows model accuracy under different values of $s$.
-The bar plot indicates the average number of hyperedges in the sampled subgraph, reflecting its exploration range.
-Figure 6: Training time vs. accuracy of random walk and $s$-walk. Bars indicate training time, markers denote accuracy.
+We investigate the impact of three key hyperparameters: the overlap size $s$, the subgraph sampling ratio $r$, and the walk length $l$ (see Appendix D.3).
+[…]
 Overlap size $s$.
-Figure 5 illustrates how model performance varies with different values of $s$, where model performance first improves and then degrades as $s$ increases.
-This trend can be attributed to the trade-off between structural coherence and exploration: when the exploration range remains sufficiently broad, a larger $s$ promotes more cohesive subgraphs while preserving essential contextual semantics.
-However, excessively strict overlap constraints (e.g., ss=10) limit structural diversity and impair performance.
-Figure 6 further indicates that increasing $s$ leads to smaller subgraphs, thereby accelerating training.
-These results highlight the importance of selecting an appropriate $s$ to balance accuracy and efficiency.
+Figure 4 illustrates how model performance varies with different $s$ while fixing $l=4$.
+Model accuracy first improves and then degrades as $s$ increases, reflecting a trade-off between structural coherence and exploration range.
+Moderate overlap constraints (e.g., $s=3$) encourage cohesive subgraphs while preserving contextual diversity, leading to the best performance in most cases.
+In contrast, $s=1$ reduces to a vanilla random walk with weak structural correlation, while overly large overlap constraints (e.g., $s=10$) greatly restrict candidate hyperedges, reducing contextual diversity and impairing performance.
+Subgraph sampling ratio $r$.
+Figure 5: Training time and accuracy under different sampling ratios $r$. $\Delta t$ indicates the relative speedup.
+[…]
 5 Related Work
-Our work is closely related to self-supervised learning on hypergraphs and representation learning on text-attributed graphs (see Appendix C.1).
+Our work is closely related to self-supervised learning on hypergraphs and representation learning on text-attributed graphs (see Appendix E.1).
 5.1 Self-supervised Learning on Hypergraphs
-SSL provides a promising alternative for hypergraph representation learning without expensive labels.
-Most existing HSSL methods focus on contrastive learning.
-HyperGCL (Wei et al.
-2022) introduces a learnable augmentation strategy, while TriCL (Lee and Shin 2023) and SE-HSSL (Li et al.
-2024) design hierarchical objectives across multiple granularities.
-Beyond contrastive approaches, generative SSL has also emerged, with HypeBoy (Kim et al.
-2024a) proposing a hyperedge-filling task.
-Other efforts address different settings, such as VilLain (Lee, Lee, and Shin 2024), which targets feature-absent scenarios.
-While effective in modeling high-order structure, these methods overlook the rich textual semantics in TAHGs.
-HyperBERT (Bazaga, Lio, and Micklem 2024) pioneers SSL on TAHGs by jointly encoding textual semantics and hypergraph structure, but it incurs high computational overhead and suffers from poor scalability.
-To bridge the gap, we explore a scalable and effective contrastive learning framework for TAHGs.
+SSL has emerged as a promising paradigm for hypergraph representation learning without expensive labels.
+Most existing HSSL methods are based on contrastive learning.
+HyperGCL Wei et al.
+(2022) introduces a learnable augmentation strategy, while TriCL Lee and Shin (2023) and SE-HSSL Li et al.
+(2024) design hierarchical contrastive objectives across multiple granularities.
+Beyond contrastive methods, generative SSL has also been explored, with HypeBoy Kim et al.
+(2024a) proposing a hyperedge-filling task.
+Other efforts explore alternative settings, such as feature-absent scenarios in VilLain Lee et al.
+(2024).
+While effective for modeling hypergraphs, these methods largely overlook the rich textual semantics in TAHGs, resulting in suboptimal performance.
+HyperBERT Bazaga et al.
+(2024) pioneers SSL on TAHGs by jointly optimizing text and hypergraph encoders through contrastive learning.
+However, its contrastive signals are limited to 1-hop hyperedge relations, restricting the modeling of long-range dependencies.
 6 Conclusion
-We propose HiTeC, a hierarchical contrastive learning framework with semantic-aware enhancement, to address limitations of existing HCL methods in TAHGs.
-To ensure scalability, HiTeC adopts a two-stage decoupled paradigm.
-The first stage pre-trains a text encoder with a structure-aware contrastive objective, alleviating the graph-agnostic nature of conventional encoders.
-In the second stage, two semantic-aware augmentation strategies, including prompt-enhanced text augmentation and semantic-aware hyperedge drop, are designed to generate more informative hypergraph views.
-Furthermore, we introduce a hierarchical contrastive objective, where a novel $s$-walk-based subgraph-level contrast captures long-range structural dependencies beyond node and hyperedge scopes.
-Extensive experiments on six real-world datasets and multiple downstream tasks demonstrate the effectiveness and scalability of HiTeC.
+We propose HiTeC, an effective contrastive learning framework tailored for TAHGs.
+HiTeC adopts a two-stage paradigm that first injects topological information into text representations and then performs semantic-aware hierarchical contrastive learning on TAHGs.
+Specifically, two semantic-aware augmentation strategies are introduced to reduce perturbation noise, while an $s$-walk-based subgraph-level contrast captures long-range dependencies beyond node- and hyperedge-level scopes.
+Extensive experiments on six real-world datasets and multiple downstream tasks demonstrate the effectiveness and robustness of HiTeC.
+Limitations
+There are two limitations in our work: (1) The construction of the structure-contextualized text view relies on a handcrafted template.
+Although the proposed template demonstrates strong empirical effectiveness across multiple datasets, its optimal form may vary across different domains and tasks.
+Future work may explore more adaptive text-level augmentation.
+(2) The structure-aware pre-training stage introduces additional training overhead compared with conventional HCL methods.
+Nevertheless, HiTeC remains more efficient than prior TAHG methods such as HyperBERT due to its decoupled training paradigm (see Appendix D.5).
 Appendix
-Appendix A Additional Experimental Setup
-A.1 Datasets
+Appendix A Additional Template Details
+Component Description Template
+[DK] Domain Knowledge You are in an Academic / E-commerce / … hypergraph.
+Each node represents a Paper / Product / …, and hyperedges capture Co-citation / Co-purchasing / … relationships among groups of nodes.
+[GT] Global Topology Your node degree is value, indicating a Sparse / Dense connectivity pattern. You participate in the following hyperedges: Hyperedge IDs.
+$t$ Raw Text Your title is Title. Your abstract or description is Abstract / Description.
+[NC] Neighbor Context Below are brief descriptions of your neighboring nodes.
+For neighbor 1, the title is: Neighbor 1 Title and the abstract or description is :Neighbor 1 Abstract / Description.
+For neighbor 2 …
+Table 3: Structure-contextualized text view template.
+As shown in Table 3, the structure-contextualized view follows a fixed template that concatenates contextual components into a single sequence: $\tilde{t}_{1,i}=[\texttt{[DK]}\,||\,\texttt{[GT]}\,||\,\texttt{[NC]}\,||\,t_{i}].$ All components are deterministically constructed from the hypergraph structure and node attributes, without external knowledge.
+The number of neighboring nodes in [NC] is limited to 10 to control sequence length and ensure stable training.
+This design enriches structural and semantic context while preserving semantic consistency.
+Appendix B Complexity Analysis
+We analyze the computational complexity of the forward phase of HiTeC during training.
+Text encoder pretraining.
+Let $l_{p}$ denote the number of layers in the PLM, $\bar{L}$ the average token length, and $d$ the hidden dimension.
+The PLM encoding costs $\mathcal{O}\!\left(l_{p}|\mathcal{V}|(\bar{L}d^{2}+\bar{L}^{2}d)\right)$.
+The triplet margin objective costs $\mathcal{O}(|\mathcal{V}|\,d)$.
+Therefore, the overall complexity is $\mathcal{O}\!\left(l_{p}|\mathcal{V}|(\bar{L}d^{2}+\bar{L}^{2}d)+|\mathcal{V}|d\right)$.
+Hypergraph encoder pretraining.
+Let $l_{h}$ be the number of HNN layers and $\delta_{e}$ the average hyperedge size.
+The hypergraph propagation costs $\mathcal{O}\!\left(l_{h}(|\mathcal{E}|\delta_{e}d+|\mathcal{V}|d^{2})\right)$.
+For contrastive learning with $m$ sampled negatives (a constant), the node-, hyperedge-, and subgraph-level objectives incur costs of $\mathcal{O}(m|\mathcal{V}|d)$, $\mathcal{O}(m|\mathcal{E}|d)$, and $\mathcal{O}(mKd)$, respectively, where $K$ denotes the number of sampled subgraphs.
+Thus, the total complexity is $\mathcal{O}(l_{h}(|\mathcal{E}|\delta_{e}d+|\mathcal{V}|d^{2})+m(|\mathcal{V}|+|\mathcal{E}|+K)d)$.
+Appendix C Additional Experimental Setup
+C.1 Datasets
 Since no publicly available benchmark datasets are specifically designed for Text-Attributed Hypergraphs (TAHGs), we construct six datasets spanning two domains: academic co-citation networks and E-commerce platforms.
@@
 Co-citation hypergraphs.
-We construct two academic hypergraphs, Citeseer (Giles, Bollacker, and Lawrence 1998) and Cora (Sen et al.
-2008), based on the standard citation networks, where each node represents a scientific publication.
-The textual attributes of each node, including the title and abstract, are directly adopted from prior works (He et al.
-2024; Wang et al.
-2024).
-To capture high-order relationships, we define a hyperedge as the set of all papers that co-cite the same reference.
+We construct two academic hypergraphs, Citeseer Giles et al.
+(1998) and Cora Sen et al.
+(2008), from standard citation networks, where each node represents a scientific publication.
+Node textual attributes, including titles and abstracts, are adopted from prior works He et al.
+(2024); Wang et al.
+(2024).
+To capture high-order relations, we define each hyperedge as the set of papers co-citing the same reference.
 E-commerce hypergraphs.
-We further construct four e-commerce hypergraphs from the Amazon product network (Ni, Li, and McAuley 2019): Book-History, Ele-Photo, Ele-Computers, and Sports-Fitness.
-In these datasets, each node represents a product, and edges reflect frequent co-purchase or co-browsing behavior.
-The raw textual attributes of nodes are obtained from the annotations provided by (Yan et al.
-2023).
-Specifically, in Book-History, textual features include book titles and descriptions; in Ele-Computers and Ele-Photo, they consist of high-rated reviews and product summaries; in Sports-Fitness, only product titles are used.
-To induce high-order structures, we adopt a max-clique-based reconstruction strategy (Bron and Kerbosch 1973; Wang and Kleinberg 2024), where hyperedges are formed by identifying maximal cliques in the co-purchase graph.
-A.2 Baselines
-We compare our method against seven strong self-supervised learning (SSL) baselines, covering both graph-based and hypergraph-based paradigms.
+We further construct four e-commerce hypergraphs from the Amazon product network Ni et al.
+(2019): Book-History, Ele-Photo, Ele-Computers, and Sports-Fitness.
+Each node represents a product, while edges reflect frequent co-purchase or co-browsing behavior.
+Node textual attributes are adopted from Yan et al.
+(2023).
+Specifically, Book-History uses book titles and descriptions; Ele-Computers and Ele-Photo use high-rated reviews and product summaries; Sports-Fitness only uses product titles.
+To induce high-order structures, we adopt a max-clique-based reconstruction strategy Bron and Kerbosch (1973); Wang and Kleinberg (2024), where hyperedges are formed from maximal cliques in the original graph.
+C.2 Baselines
+We compare our method against eight representative self-supervised learning (SSL) baselines, covering graph-based, hypergraph-based, and text-attributed hypergraph (TAHG)-based paradigms.
 Graph-based SSL.
-In this work, we adopt three widely used graph-based SSL frameworks: two contrastive learning methods, GraphCL (You et al.
-2020) and BGRL (Thakoor et al.
-2022), and one generative method, GraphMAE (Hou et al.
-2022).
-To adapt these methods to hypergraphs, we apply clique expansion following the protocol in (Lee and Shin 2023).
-For textual feature extraction, we employ shallow encoders (SE) implemented via Skip-Gram (Mikolov et al.
-2013), as well as GIANT (Chien et al.
-2022a), a graph-agnostic text encoder designed for raw textual attributes.
+In this work, we adopt three widely used graph-based SSL frameworks: two contrastive learning methods, GraphCL You et al.
+(2020) and BGRL Thakoor et al.
+(2022), and one generative method, GraphMAE Hou et al.
+(2022).
+To adapt these methods to hypergraphs, we apply clique expansion following the protocol in Lee and Shin (2023).
+For textual feature extraction, we employ shallow encoders (SE) implemented via Skip-Gram Mikolov et al.
+(2013), as well as GIANT Chien et al.
+(2022a), a graph-agnostic text encoder designed for raw textual attributes.
 Hypergraph-based SSL.
-We include four state-of-the-art hypergraph SSL methods: two contrastive approaches, TriCL (Lee and Shin 2023) and SE-HSSL (Li et al.
-2024); one generative method, HypeBoy (Kim et al.
-2024a); and VilLain (Lee, Lee, and Shin 2024), which targets the feature-absent setting.
-All methods adopt the same SE setup and are additionally evaluated with two pre-trained language models: BERT-Base (Devlin et al.
-2019) and RoBERTa-Base (Liu et al.
-2019).
-A.3 Evaluation Protocol
+We consider four state-of-the-art self-supervised learning methods specifically designed for hypergraphs.
+These include two contrastive approaches, TriCL Lee and Shin (2023) and SE-HSSL Li et al.
+(2024); one generative framework, HypeBoy Kim et al.
+(2024a); and VilLain Lee et al.
+(2024), which is tailored for feature-absent hypergraph settings.
+For a fair comparison, all methods adopt the same shallow encoder (SE).
+In addition, we further evaluate these baselines by replacing the SE with two pre-trained language models, BERT-Base Devlin et al.
+(2019) and RoBERTa-Base Liu et al.
+(2019).
+TAHG-based SSL.
+We further include HyperBERT Bazaga et al.
+(2024), to the best of our knowledge, the only existing SSL method specifically designed for text-attributed hypergraphs.
+HyperBERT jointly models textual semantics and hypergraph structure using a Transformer-based architecture, in which hypergraph-aware layers are introduced to integrate structural signals with text representations.
+However, this joint modeling paradigm incurs substantial computational and memory overhead, limiting its scalability to large hypergraphs, as validated in our experiments.
+C.3 Evaluation Protocol
 We evaluate our method on two standard hypergraph learning tasks: node classification and hyperedge prediction.
-After pretraining with HiTeC, we adopt a linear evaluation protocol (Lee and Shin 2023; Li et al.
-2024; Kim et al.
-2024a), where the pretrained node embeddings are frozen and used as input features for downstream classifiers.
+After pretraining with HiTeC, we adopt a linear evaluation protocol Lee and Shin (2023); Li et al.
+(2024); Kim et al.
+(2024a), where the pretrained node embeddings are frozen and used as input features for downstream classifiers.
 Node classification.
-Following (Lee and Shin 2023; Li et al.
-2024), we randomly split each dataset into 10%/10%/80% for training, validation, and testing.
+Following Lee and Shin (2023); Li et al.
+(2024), we randomly split each dataset into 10%/10%/80% for training, validation, and testing.
 Then, a simple linear classifier is trained on top of the fixed node embeddings using $\ell_{2}$-regularized logistic regression.
@@
 Hyperedge prediction.
-Following (Kim et al.
-2024a), we need to generate negative hyperedge samples to evaluate a model on the hyperedge prediction task.
-In our experiments, we adopt Clique Negative Sampling (CNS) (Yu et al.
-2025), a widely used strategy for constructing negative hyperedge samples.
+Following Kim et al.
+(2024a), we need to generate negative hyperedge samples to evaluate a model on the hyperedge prediction task.
+In our experiments, we adopt Clique Negative Sampling (CNS) Yu et al.
+(2025), a widely used strategy for constructing negative hyperedge samples.
 Specifically, for a given hyperedge $e$, we replace a node $u\in e$ with a node $v\notin e$ such that $v$ is connected to all other nodes in $e$, i.e., $(e\setminus\{u\})\cup\{v\}$.
@@
 Same as the node classification task, we report the average test accuracy and standard deviation over 20 random splits, each with 5 initializations.
-A.4 Implementation Details
-All experiments are conducted on a single NVIDIA A5000 GPU (24GB memory).
+C.4 Implementation Details
+All experiments are conducted on a single NVIDIA A800 PCIe GPU with 80GB of memory.
 For all baselines, we adopt their official implementations with default settings.
-For HiTeC, we adopt BERT-Base (Devlin et al.
-2019) as the text encoder and apply LoRA (Hu et al.
-2022) for parameter-efficient fine-tuning.
+For HiTeC, we adopt BERT-Base Devlin et al.
+(2019) as the default text encoder and apply LoRA Hu et al.
+(2022) to the query and value matrices of the self-attention layers for parameter-efficient fine-tuning, while keeping the PLM backbone frozen.
+Unless otherwise specified, the LoRA configuration is set to $r=16$, $\alpha=32$, and dropout =0.05=0.05.
+We additionally evaluate other text encoder backbones, including DistilBERT Sanh et al.
+(2019), RoBERTa-Base Liu et al.
+(2019), and DeBERTa-Base He et al.
+(2020).
 The text encoder is trained for 5 epochs on small datasets (Citeseer, Cora), 4 epochs on medium datasets (History, Photo), and 3 epochs on large datasets (Computers, Fitness), using a learning rate of 2e-5.
 The hypergraph encoder is implemented as a single-layer module.
-We tune the contrastive loss weight $\lambda_{e}$ and $\lambda_{s}$ over $\{1,2,3,4\}$.
-The overlap threshold $s$ and subgraph sampling ratio $r\%$ are tuned from {1, 2, 3, 4, 5, 10} and {10%, 20%, 30%, 40%, 50%}, respectively.
-To ensure a fair comparison, we tune the learning rate from {0.01, 0.001, 1e-4, 5e-4} and weight decay from {0, 1e-5} for all methods.
-Figure 7: Training time and accuracy under different sampling ratios $r\%$.
-Bar plots represent training time (in minutes), while line plots indicate the corresponding accuracy.
-We highlight the relative speedup $\Delta t$ of $r\%{=}30\%$ over $r\%{=}50\%$.
-All main results are reported with $r\%{=}30\%$.
+We tune the contrastive loss weights $\lambda_{e}$ and $\lambda_{s}$ over $\{1,2,3,4\}$.
+The overlap threshold $s$ and walk length $l$ are tuned from $\{1,2,3,4,5,10\}$, the subgraph sampling ratio $r$ in $\{10,20,30,40,50\}$.
+To ensure a fair comparison, we tune the learning rate from $\{0.01,0.001,1e\text{-}4,5e\text{-}4\}$ and weight decay from $\{0,1e\text{-}5\}$ for all methods.
+Appendix D Additional Experiment Results
 Method Citeseer Cora History Photo Computers Fitness
-Graph SSL
-GraphCL + SE 69.48 ± 2.07 59.05 ± 2.16 65.65 ± 1.06 67.48 ± 1.18 68.66 ± 1.33 61.22 ± 1.34
+Graph SSL GraphCL + SE 69.48 ± 2.07 59.05 ± 2.16 65.65 ± 1.06 67.48 ± 1.18 68.66 ± 1.33 61.22 ± 1.34
 GraphCL + GIANT 72.70 ± 2.06 55.48 ± 1.96 63.02 ± 1.61 68.78 ± 1.16 70.87 ± 0.87 63.02 ± 0.01
@@
 GraphMAE + GIANT 72.25 ± 1.62 66.30 ± 2.19 64.19 ± 0.23 64.83 ± 0.20 66.10 ± 1.20 59.21 ± 0.88
-HSSL
-TriCL + SE 68.06 ± 1.75 62.43 ± 2.41 67.18 ± 0.24 69.21 ± 0.19 70.63 ± 0.14 63.87 ± 0.30
+HSSL TriCL + SE 68.06 ± 1.75 62.43 ± 2.41 67.18 ± 0.24 69.21 ± 0.19 70.63 ± 0.14 63.87 ± 0.30
 TriCL + BERT 71.36 ± 1.78 64.02 ± 2.33 70.18 ± 0.31 71.71 ± 0.21 73.20 ± 0.17 65.49 ± 0.40
@@
 HypeBoy + RoBERTa 73.47 ± 2.10 66.51 ± 1.87 OOM OOM OOM OOM
-VilLain (w/o feat.) 74.25 ± 1.90 63.25 ± 2.57 63.18 ± 0.35 65.02 ± 0.23 65.85 ± 0.28 OOM
-HyperBERT 51.84 ± 2.03 45.32 ± 1.47 OOM OOM OOM OOM
-HiTeC 74.60 ± 1.84 65.91 ± 2.32 71.83 ± 0.21 72.56 ± 0.17 73.91 ± 0.14 67.32 ± 0.09
-Table 3: Evaluation results on hyperedge prediction (% mean ± std).
+VilLain (w/o feat.) 74.25 ± 1.90 63.25 ± 2.57 63.18 ± 0.35 65.02 ± 0.23 65.85 ± 0.28 62.92 ± 0.26
+HyperBERT 71.84 ± 2.03 64.48 ± 2.10 69.52 ± 0.20 69.34 ± 0.18 70.40 ± 0.15 OOM
+HiTeC 75.22 ± 1.87 65.91 ± 2.32 71.83 ± 0.21 72.56 ± 0.17 73.91 ± 0.14 67.32 ± 0.09
+Table 4: Evaluation results on hyperedge prediction task(% mean $\pm$ std).
 The best and second-best results are highlighted in bold and underline, respectively.
-‘w/o feat.” means without node features.
-OOM denotes out-of-memory on a 24GB GPU.
-Dataset DistilBERT BERT RoBERTa DeBERTa Citeseer 67.54 ± 1.49 66.38 ± 1.73 65.09 ± 1.26 64.83 ± 4.65 Cora 74.55 ± 0.02 74.07 ± 0.88 72.32 ± 0.91 70.41 ± 2.24 History 78.05 ± 0.18 79.81 ± 0.16 74.77 ± 0.24 76.20 ± 0.12 Photo 56.90 ± 0.31 59.64 ± 0.16 54.15 ± 0.40 52.54 ± 0.08 Computers 50.64 ± 0.25 55.81 ± 0.12 52.36 ± 0.21 51.48 ± 0.56 Fitness 62.97 ± 0.08 67.18 ± 0.19 64.98 ± 0.78 65.29 ± 0.14
-Table 4: Different text encoder backbones.
-Appendix B Additional Experiment Results
-B.1 Hyperedge Prediction Evaluation
+“w/o feat.” denotes methods without node features.
+OOM indicates that the method runs out of memory on a single 80GB GPU.
+D.1 Hyperedge Prediction Evaluation
 We further evaluate all models on the hyperedge prediction task, with results presented in Table 4.
@@
 Overall, HiTeC exceeds the second-best method by 1.65% and 1.15% on History and Fitness, respectively, demonstrating its robustness and scalability across diverse hypergraph scenarios.
-B.2 Impact of Text Encoder Backbones
-We conduct an ablation study on different text encoder backbones, including DistilBERT (Sanh et al.
-2019), BERT-Base (Devlin et al.
-2019), RoBERTa-Base (Liu et al.
-2019), and DeBERTa-Base (He et al.
-2020), as shown in Table 4.
-Among these, BERT achieves the highest overall performance, followed by DistilBERT.
-These results suggest that selecting a suitable text encoder for each dataset is important.
-Nevertheless, our framework is compatible with various backbones.
-In our main experiments, we adopt BERT as the default text encoder backbone.
-B.3 Parameter Sensitivity Analysis of Subgraph Sampling Ratio $r\%$
+D.2 Impact of Text Encoder Backbones
+We evaluate different text encoder backbones, including DistilBERT Sanh et al.
+(2019), BERT-Base Devlin et al.
+(2019), RoBERTa-Base Liu et al.
+(2019), and DeBERTa-Base He et al.
+(2020), as shown in Table 5.
+Among them, BERT-Base achieves the best overall performance, followed by DistilBERT.
+The results indicate that backbone selection can noticeably affect performance across datasets.
+Nevertheless, our framework remains compatible with diverse text encoders.
+Based on the overall results, we adopt BERT as the default backbone in the main experiments.
+Dataset DistilBERT BERT RoBERTa DeBERTa Citeseer 67.54 ± 1.49 66.38 ± 1.73 65.09 ± 1.26 64.83 ± 4.65 Cora 74.55 ± 0.02 74.07 ± 0.88 72.32 ± 0.91 70.41 ± 2.24 History 78.05 ± 0.18 79.81 ± 0.16 74.77 ± 0.24 76.20 ± 0.12 Photo 56.90 ± 0.31 59.64 ± 0.16 54.15 ± 0.40 52.54 ± 0.08 Computers 50.64 ± 0.25 55.81 ± 0.12 52.36 ± 0.21 51.48 ± 0.56 Fitness 62.97 ± 0.08 67.18 ± 0.19 64.98 ± 0.78 65.29 ± 0.14
+Table 5: Different text encoder backbones.
+D.3 Parameter Sensitivity Analysis of $s$-walk length $l$
+Figure 6: Sensitivity analysis of walk length $l$.
+To investigate the effect of walk length on HiTeC, we report the results under different $l\in\{1,2,3,4,5,10\}$ while fixing $s=3$.
+As shown in Figure 6, model performance generally first improves and then degrades as $l$ increases.
+This trend suggests that moderate walk lengths help capture richer high-order structural dependencies by exploring broader hypergraph contexts.
+In contrast, short walks (e.g., $l=1$) provide limited structural context, while excessively long walks (e.g., $l=10$) may introduce noisy or weakly related hyperedges, both leading to suboptimal performance.
+Overall, $l=4$ achieves a favorable balance between structural exploration and semantic consistency.
+D.4 Parameter Sensitivity Analysis of Subgraph Sampling Ratio $r$
 […]
-Appendix C Additional Related Work
-C.1 Representation Learning on TAGs
+Figure 7: Training time and accuracy under different sampling ratios $r$.
+Bar plots report training time (in seconds), while line plots show the corresponding accuracy.
+We highlight the relative speedup $\Delta t$ of $r=30$ over $r=50$.
+$s$ and $l$ are fixed to 3 and 4, respectively.
+All main results are selected with $r=30$.
+D.5 Efficiency Evaluation
+We evaluate training time to assess the efficiency of HiTeC.
+Specifically, we first compare the text encoder pre-training time between HiTeC and HyperBERT, and then compare the hypergraph encoder pre-training time of HiTeC with two representative HCL methods (TriCL and SE-HSSL).
+[…]
+Method Citeseer Cora History Photo Computers HyperBERT 4h 30m 5h 22m 29h 42m 33h 40m 56h 04m HiTeC 2h 22m 2h 56m 18h 31m 19h 55m 28h 44m
+Table 6: Comparison of text encoder pretraining time between HyperBERT and HiTeC.
+Hypergraph encoder pretraining time evaluation.
+We further compare the hypergraph encoder pre-training time of HiTeC with two representative hypergraph-based HCL methods.
+As shown in Figure 8, training time is plotted on a logarithmic scale for clearer visualization, where HiTeC achieves performance comparable to SE-HSSL while being at least $3\times$ faster than TriCL on large-scale datasets.
+This efficiency gain mainly stems from two factors: (1) The subgraph-level contrastive objective operates on a small subset of sampled subgraphs (see Appendix D.4); and (2) The preprocessing steps are executed once and reused across multiple training runs.
+Overall, HiTeC strikes a favorable trade-off between training efficiency and accuracy.
+Figure 8: Comparison of hypergraph encoder pretraining time among HiTeC, SE-HSSL, and TriCL. log(·) denotes the natural logarithm.
+Appendix E Additional Related Work
+E.1 Representation Learning on TAGs
 Since hypergraphs can be transformed into graphs via clique expansion, our work is closely related to representation learning on text-attributed graphs (TAGs).
-Recent TAG methods leverage pre-trained language models to generate numerical features, typically in a cascade (He et al.
-2024; Wang et al.
-2024) or nested (Yang et al.
-2021; Zhao et al.
-2023) manner.
+Recent TAG methods leverage pre-trained language models to generate numerical features, typically in a cascade He et al.
+(2024); Wang et al.
+(2024) or nested Yang et al.
+(2021); Zhao et al.
+(2023) manner.
 While most existing TAG methods are supervised, emerging efforts explore self-supervised learning (SSL) to reduce annotation costs.
-GIANT (Chien et al.
-2022a) makes an early attempt by aligning textual and structural features via an extreme multi-label classification objective.
-More recently, GAugLLM (Fang et al.
-2024) proposes an augmentation framework that utilizes large language models (LLMs) to perturb both node features and graph structure, while HASH-CODE (Zhang et al.
-2024) formulates a multi-objective contrastive learning strategy tailored for TAGs.
-However, adapting these methods to hypergraphs via clique expansion incurs substantial structural loss, limiting their effectiveness on TAHGs and motivating the need for dedicated HSSL approaches.
+GIANT Chien et al.
+(2022a) makes an early attempt by aligning textual and structural features via an extreme multi-label classification objective.
+More recently, GAugLLM Fang et al.
+(2024) proposes an augmentation framework that utilizes large language models (LLMs) to perturb both node features and graph structure, while HASH-CODE Zhang et al.
+(2024) formulates a multi-objective contrastive learning strategy tailored for TAGs.
+However, hypergraphs fundamentally differ from ordinary graphs in their high-order structural modeling.
+Converting hypergraphs into ordinary graphs via clique expansion Bron and Kerbosch (1973) inevitably incurs substantial structural information loss Lee and Shin (2023); Li et al.
+(2024), limiting their effectiveness on TAHGs and motivating the need for dedicated HSSL approaches.